### v2

# Revisiting Schwarzschild black hole singularity through string theory

Abstract
In this letter, we derive the singular condition for black holes and demonstrate the potential resolution of the black hole singularity in general relativity using non-perturbative $\alpha^{\prime}$ corrections of string theory. This work is motivated by the Belinskii, Khalatnikov and Lifshitz (BKL) proposal, which suggests that the structure of the black hole interior in vacuum Einstein’s equations can be transformed into the Kasner universe near the singularity. This transformation allows for the description of the black hole interior using the $O\left(d,d\right)$ invariant anisotropic Hohm-Zwiebach action, which includes all orders of $\alpha^{\prime}$ corrections.
preprint: APS/123-QED
Black holes are the most important objects in our universe, as understanding their internal structure and microstate could reveal the fundamental nature of physics. However, gravitational collapse inevitably leads to a curvature singularity at the center of a black hole, a feature that remains unavoidable in the framework of Einstein’s gravity [1, 2]. This suggests that general relativity is incomplete and breaks down in the vicinity of this singularity. String theory, a candidate theory of quantum gravity, is expected to modify the general relativity by including higher-derivative corrections and resolving this longstanding issue. However, to date, it has only provided insights into two-dimensional string black holes [3, 4], leaving unresolved the singularity problem of the four-dimensional black hole in general relativity. This unbridgeable gap arises from the unknown non-perturbative aspects of string theory when approaching the curvature singularity.
In this letter, we aim to investigate whether the singularity of the Schwarzschild and rotating, accreting black holes can be resolved by higher-derivative $\alpha^{\prime}$ corrections of string theory. This work is motivated by the recent progress in classifying all orders of $\alpha^{\prime}$ corrections for specific backgrounds by Hohm and Zwiebach [5, 6, 7]. In previous work, Sen has demonstrated that when the classical fields, including the metric, Kalb-Ramond field and dilaton, are independent of $m$ coordinates, the field theory exhibits an $O\left(m,m\right)$ symmetry to all orders in $\alpha^{\prime}$ [8, 9]. This result has been verified for a few orders in $\alpha^{\prime}$, indicating that the standard form of $O\left(d,d\right)$ transformations can be preserved through suitable field redefinitions, even in the presence of $\alpha^{\prime}$ corrections [10, 8, 9, 11, 12, 13]. Based on these results, Hohm and Zwiebach conjectured that the standard form of $O\left(d,d\right)$ transformations also preserves for all orders in $\alpha^{\prime}$, enabling the classification of all $O\left(d,d\right)$ invariant $\alpha^{\prime}$ corrections. Since the corresponding $O\left(d,d\right)$ invariant action involves only first-order derivatives, it can be exactly solved. Consequently, it raises the question of whether black hole singularities can be cured based on this new development. Recent studies have successfully resolved the big bang singularity [14, 15, 16, 17, 18, 19] and the curvature singularities of two-dimensional string black holes [20, 21, 22, 23] in this framework. However, this approach still cannot resolve the singularity of the Schwarzschild and rotating, accreting black holes. The challenge arises from the fact that the spherical part of the black hole metric does not exhibit $O\left(d,d\right)$ symmetry. Therefore, directly applying the Hohm-Zwiebach action becomes impossible.
Encouragingly, the general structure near the black hole singularity, namely Kasner universe, exhibits $O\left(d,d\right)$ symmetry, opening avenues for understanding the black hole singularity in non-perturbative string theory. To be specific, let us recall the tree-level string effective action:
$I_{\mathrm{String}}=\frac{1}{16\pi G_{D}}\int d^{D}x\sqrt{-g}e^{-2\phi}\left(R+4\left(\partial\phi\right)^{2}\right).$ (1)
where we adopt $116\pi G_{D}=$ 1, $\phi\left(x\right)$ denotes a physical dilaton and we set Kalb-Ramond field $b_{\mu\nu}=0$ for simplicity. Moreover, when $\phi$ is a constant, the action reduces to the Einstein-Hilbert action and we will consider the stringy corrections to this action in this letter. On the other hand, since the spacetime dimension of bosonic string theory is fixed to be $26$, we can see the manifold as a direct product of the black hole and a compact internal space [24]. Due to the fact that the decoupling of the black hole and internal space hold for all order of $\alpha^{\prime}$ corrections [25], we only need to consider the black hole part in the following calculations. To obtain the Kasner universe near the black hole singularity, we first consider the $D=n+3$ dimensional Schwarzschild-Tangherlini black hole [26, 27]:
$ds^{2}=-\left(1-\left(\frac{r_{0}}{r}\right)^{n}\right)dt^{2}+\frac{dr^{2}}{1-\left(\frac{r_{0}}{r}\right)^{n}}+r^{2}d\Omega_{n+1}^{2},$ (2)
where $d\Omega_{n+1}^{2}$ denotes the metric of an $\left(n+1\right)$ dimensional sphere, and $r_{0}$ is an event horizon. Inside the event horizon, the metric can be transformed into the Kasner metric near the singularity [28]:
$ds^{2}=-d\tau^{2}+\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}\tau^{2\beta_{i}}dx_{i}^{2}.$ (3)
The spacelike singularity is now located at $\tau=0$ which corresponds to $r=0$ of Schwarzschild black hole. Moreover, in order to satisfy the vacuum Einstein’s equations, the Kasner exponents $\beta_{i}$ must satisfy the following constraints
$\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}\beta_{i}^{2}=1,\qquad\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}\beta_{i}=1.$ (4)
For the four-dimensional Schwarzschild black hole, we have $\left(\beta_{1},\beta_{2},\beta_{3}\right)=\left(-\frac{1}{3},\frac{2}{3},\frac{2}{3}\right)$. One may wonder whether the initial spherical symmetry is broken in the Kasner metric (3). Near the singularity, the curvature becomes large enough that the evolution of the black hole geometry at different spatial points decouples. It means that the partial differential Einstein’s equations become ordinary differential equations (ODEs) with respect to time [29]. Therefore, there is no reason to assume the initial spherical symmetry is also maintained in ODEs’ solution. The BKL approximation (with the Kasner metric as its solution) is sufficient to describe this region. It depicts an asymmetrically spherical shell chaotically collapsing near the singularity.
Apart from the Schwarzschild black hole, the Kerr black hole is highly interesting. However, the presence of an inner horizon in a rotating black hole transforms the spacelike singularity into a timelike singularity. Poisson and Israel studied the mass inflation instability at the inner horizon, which leads to the formation of a null singularity at the inner horizon [30]. In subsequent work, Burko found that if the radiation power law does not drop off quickly, a spacelike singularity rather than a null singularity will be formed. In other words, the inner horizon can be superseded by the BKL singularity [31]. Hamilton then numerically studied the evolution of the inner horizon of a rotating, accreting black hole, which implies that the spacetime almost undergoes a BKL collapse [32]. Subsequently, the analytic model of this collapse, called the inflationary Kasner solution, was constructed, including two Kasner epochs: an inflation epoch with Kasner exponents $\left(\beta_{1},\beta_{2},\beta_{3}\right)=\left(1,0,0\right)$ and a collapse epoch characterized by Kasner exponents $\left(\beta_{1},\beta_{2},\beta_{3}\right)=\left(-\frac{1}{3},\frac{2}{3},\frac{2}{3}\right)$ approaching the spacelike singularity, which approximates a Schwarzschild geometry [33].
Since the Kasner universe only depends on $\tau$, the string action in this background possesses the $O\left(d,d\right)$ symmetry. This allows us to study the effects of string theory non-perturbatively in near singularity region and examine how it may potentially resolve the black hole singularities (the previous discussions on anisotropic Hohm-Zwiebach action can be found in references [34, 35, 17, 36]. Although the spacelike singularity of isotropic string cosmology has been resolved in many references [37, 38], the anisotropic case is highly non-trivial and lacks any progress on its exact solution [36]). Thus, let us use the following ansatz:
$ds^{2}$ $=$ $-d\tau^{2}+g_{ij}dx^{i}dx^{j},$
$g_{ij}$ $=$ $\mathrm{diag}\left(a_{1}\left(\tau\right)^{2},a_{2}\left(\tau\right)^{2},\ldots,a_{i}\left(\tau\right)^{2}\right),$ (5)
where $i=1,\ldots,d$. The corresponding anisotropic Hohm-Zwiebach action is given by:
$I_{\mathrm{HZ}}$ $=$ $\int d^{D}x\sqrt{-g}e^{-2\phi}\left(R+4\left(\partial\phi\right)^{2}\right.$ (6)
$\left.+\frac{1}{4}\alpha^{\prime}\left(R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}+\cdots\right)+\alpha^{\prime 2}\left(\cdots\right)+\cdots\right)$
$=$ $\int d\tau e^{-\Phi}\left(-\dot{\Phi}^{2}-\frac{1}{8}\mathrm{Tr}\left(\dot{\mathcal{S}}^{2}\right)\right.$
$+\alpha^{\prime}c_{2,0}\mathrm{Tr}\left(\dot{\mathcal{S}}^{4}\right)+\alpha^{\prime 2}c_{3,0}\mathrm{Tr}\left(\dot{\mathcal{S}}^{6}\right)$
$+\alpha^{\prime 3}\left[c_{4,0}\mathrm{Tr}\left(\dot{\mathcal{S}}^{8}\right)+c_{4,1}\left(\mathrm{Tr}\left(\dot{\mathcal{S}}^{4}\right)\right)^{2}\right]$
[…]
where $\dot{f}\left(\tau\right)\equiv\partial_{\tau}f\left(\tau\right)$. Moreover, there is no arbitrariness for this action as a string effective action due to the form of this action is constrained by the symmetry and has removed ambiguous coefficients through the field redefinition. On the other hand, it has been verified that at order $\alpha^{\prime 3}$ of Type II string theory [39] and order $\alpha^{\prime}$ of general torus compactifications [40], they are consistent with (6) with specific coefficients $c_{i,j}$. The string vacuum is also a solution of this action [6, 16]. For a bosonic case, $c_{1,0}=-\frac{1}{8}$, $c_{2,0}=\frac{1}{64}$, $c_{3,0}=-\frac{1}{3.2^{7}}$, $c_{4,0}=\frac{1}{2^{12}}-\frac{3}{2^{12}}\zeta\left(3\right)$, $c_{4,1}=\frac{1}{2^{16}}+\frac{1}{2^{12}}\zeta\left(3\right)$ and $c_{k>4}$’s are unknown [13]. This action systematically provides all-order higher curvature corrections to Einstein’s gravity. For example, the leading order is the Einstein-Hilbert action ($\phi=\mathrm{constant}$), and the first order of $\alpha^{\prime}$ correction is the Gauss-Bonnet term. Moreover, $O\left(d,d\right)$ invariant dilaton is defined by $\Phi\equiv 2\phi-\log\sqrt{-g}$. The matrix $\mathcal{S}$ is defined as
$\mathcal{S}\equiv\left(\begin{array}[]{cc}0&g_{ij}\\ g_{ij}^{-1}&0\end{array}\right),\quad\mathrm{Tr}\left(\dot{\mathcal{S}}^{2k}\right)=\left(-4\right)^{k}2\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}^{2k},$ (7)
where $H_{i}\equiv\frac{\dot{a}_{i}\left(\tau\right)}{a_{i}\left(\tau\right)}$ denotes the Hubble parameter. In addition, the $O\left(d,d\right)$ symmetry of the action (6) is manifested by the following transformations:
$\Phi\rightarrow\Phi,\qquad a_{i}\rightarrow a_{i}^{-1},\qquad H_{i}\rightarrow-H_{i}.$ (8)
However, it is challenging to obtain the solutions of EOM from the action (6) due to the multi-trace terms. As an example, let us consider the action at the order of $\alpha^{\prime 3}$ which introduces the first multi-trace term into the action (6). The action can be written as follows:
$I_{\mathrm{HZ}}^{\left(3\right)}=\alpha^{\prime 3}\left[c_{4,0}\mathrm{Tr}\left(\dot{\mathcal{S}}^{8}\right)+c_{4,1}\left(\mathrm{Tr}\left(\dot{\mathcal{S}}^{4}\right)\right)^{2}\right].$ (9)
For the isotropic background:
$I_{\mathrm{HZ}}^{\left(3\right)}$ $=$ $\alpha^{\prime 3}\left(c_{4,0}\left(-4\right)^{4}2dH^{8}+c_{4,1}\left(-4\right)^{4}4d^{2}H^{8}\right)$ (10)
$=$ $\alpha^{\prime 3}\left(\mathrm{constant}\times H^{8}\right),$
all orders of $\alpha^{\prime}$ corrections can be straightforwardly summed in a simple manner proportional to $H^{2k}$. And it is easy to obtain the corresponding EOM. But, for the anisotropic case:
$I_{\mathrm{HZ}}^{\left(3\right)}=\alpha^{\prime 3}\left(-4\right)^{4}\left[2c_{4,0}\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}^{8}+4c_{4,1}\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}^{4}\right)^{2}\right],$ (11)
which cannot be summed together and the effects of multi-trace cannot be neglected.
The aim of this letter is to calculate the non-perturbative and non-singular black hole solution of the action (6) with all-order higher curvature corrections. To calculate its solution, we first recall the EOM of the action (6):
$\ddot{\Phi}+\frac{1}{2}\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}f_{i}\left(H_{i}\right)$ $=$ $0,$
$\left(\frac{d}{d\tau}-\dot{\Phi}\right)f_{i}\left(H_{i}\right)$ $=$ $0,$
$\dot{\Phi}^{2}+\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}g_{i}\left(H_{i}\right)$ $=$ $0,$ (12)
where
$f_{i}\left(H_{i}\right)$ $=$ $-2H_{i}-2\alpha^{\prime}H_{i}^{3}-2\alpha^{\prime 2}H_{i}^{5}$
$-8\alpha^{\prime 3}4^{4}\left[2c_{4,0}H_{i}^{7}+4c_{4,1}\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}H_{j}^{4}\right)H_{i}^{3}\right]+\cdots,$
$g_{i}\left(H_{i}\right)$ $=$ $-H_{i}^{2}-\frac{3}{2}\alpha^{\prime}H_{i}^{4}-\frac{5}{3}\alpha^{\prime 2}H_{i}^{6}$
$-7\alpha^{\prime 3}4^{4}\left[2c_{4,0}H_{i}^{8}+4c_{4,1}\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}H_{j}^{4}\right)H_{i}^{4}\right]+\cdots,$
and there is no additional constraint $\frac{d}{dH_{i}}g_{i}\left(H_{i}\right)=H_{i}\frac{d}{dH_{i}}f_{i}\left(H_{i}\right)$ for the anisotropic case. These EOM are consistent with refs. [35, 36]. When $\alpha^{\prime}=0$, the tree-level EOM (12) have a solution
$\Phi\left(\tau\right)$ $=$ $-\log\left(\tau-\tau_{0}\right)+\Phi_{0},$
$a_{i}\left(\tau\right)$ $=$ $C_{i}\left(\tau-\tau_{0}\right)^{\beta_{i}},$
$H_{i}\left(\tau\right)$ $=$ $\frac{\beta_{i}}{\tau-\tau_{0}},$ (14)
where $\tau_{0}$, $C_{i}$ and $\Phi_{0}$ are arbitrary constants. This result covers the solution of Einstein’s gravity (3) since the physical dilaton $\phi$ is a constant. The corresponding Kretschmann scalar is
$R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}=\frac{4}{\left(\tau-\tau_{0}\right)^{4}}\left(\frac{1}{2}\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}\beta_{i}^{4}-2\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}\beta_{i}^{3}+\frac{3}{2}\right).$ (15)
The spacelike singularity is located at $\tau=\tau_{0}$. Now, our aim is to remove this singularity using $\alpha^{\prime}$ corrections in string theory by the following strategy:
  • Firstly, we aim to calculate the perturbative solution of the EOM (12) in the perturbative regime as $\alpha^{\prime}\rightarrow 0$.
  • Then, in the non-perturbative regime for an arbitrary $\alpha^{\prime}$, we need to find a non-singular solution that satisfies the EOM (12) and covers arbitrary higher orders of perturbative solution with the multi-trace terms as $\alpha^{\prime}\rightarrow 0$ (the detailed calculations can be found in the Appendix).
To calculate the perturbation solution of the EOM (12), we introduce a new variable $\Omega$:
$\Omega\equiv e^{-\Phi},$ (16)
where $\dot{\Omega}=-\dot{\Phi}\Omega$ and $\ddot{\Omega}=\left(-\ddot{\Phi}+\dot{\Phi}^{2}\right)\Omega$. And the EOM (12) become
$\ddot{\Omega}-\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}h_{i}\left(H_{i}\right)\right)\Omega$ $=$ $0,$
$\frac{d}{d\tau}\left(\Omega f_{i}\left(H_{i}\right)\right)$ $=$ $0,$
$\dot{\Omega}^{2}+\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}g_{i}\left(H_{i}\right)\right)\Omega^{2}$ $=$ $0,$ (17)
where we define a new function
$h_{i}\left(H_{i}\right)\equiv\frac{1}{2}H_{i}f_{i}\left(H_{i}\right)-g\left(H_{i}\right)=\alpha^{\prime}\frac{1}{2}H_{i}^{4}+\cdots,$ (18)
Next, we assume the perturbative solutions of the EOM (17) to be:
$\Omega\left(\tau\right)$ $=$ $\Omega_{\left(0\right)}\left(\tau\right)+\alpha^{\prime}\Omega_{\left(1\right)}\left(\tau\right)+\alpha^{\prime 2}\Omega_{\left(2\right)}\left(\tau\right)+\cdots,$
$H_{i}\left(\tau\right)$ $=$ $H_{i\left(0\right)}\left(\tau\right)+\alpha^{\prime}H_{i\left(1\right)}\left(\tau\right)+\alpha^{\prime 2}H_{i\left(2\right)}\left(\tau\right)+\cdots,$
where we denote $\Omega_{i}$ and $H_{i}$ as the $i$th order of the perturbative solutions. Therefore, the perturbative solution can be calculated order by order, which is
$H_{i}\left(\tau\right)$ $=$ $\frac{\beta_{i}}{\tau-\tau_{0}}-\frac{4\beta_{i}^{3}+\beta_{i}\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}}{4\left(\tau-\tau_{0}\right)^{3}}\alpha^{\prime}+\cdots,$
$\Omega\left(\tau\right)$ $=$ $\gamma\left(\tau-\tau_{0}\right)+\frac{\gamma}{4\left(\tau-\tau_{0}\right)}\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}\beta_{i}^{4}\right)\alpha^{\prime}+\cdots,$
where $\gamma$ is an integration constant. The leading-order of (LABEL:eq:perturbed_solution) covers the geometry near the Schwarzschild singularity (3). Moreover, due to equation (16), we also have:
$\Phi\left(\tau\right)=-\log\left(\gamma\left(\tau-\tau_{0}\right)\right)-\frac{1}{4\left(\tau-\tau_{0}\right)^{2}}\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}\beta_{i}^{4}\right)\alpha^{\prime}+\cdots.$ (21)
Based on the perturbative solution (LABEL:eq:perturbed_solution) and (21), we obtain the non-perturbative and non-singular solution which satisfies the EOM (12). The dilaton solution is given by
$\Phi\left(\tau\right)=-\frac{1}{2N}\log\left(\stackrel{{\scriptstyle[}}{{k}}=0]{N}{\sum}\lambda_{k}\sigma^{2k}\right),\qquad\sigma^{2}\equiv 2d\frac{\tau^{2}}{\alpha^{\prime}},$ (22)
where $\lambda_{k}$ is fixed by the coefficients $c_{1}$, … , $c_{k}$, and the corresponding solutions for $f_{i}$, $g_{i}$ and $H_{i}$:
$f_{i}$ $=$ $-\frac{2\beta_{i}}{\sqrt{\alpha^{\prime}}}e^{\Phi\left(\tau\right)},$
$g_{i}$ $=$ $-\dot{\Phi}^{2}\stackrel{{\scriptstyle[}}{{k}}=1]{\infty}{\sum}\frac{b_{k}}{1+\sigma^{2k-2}},$
$H_{i}$ $=$ $\frac{\sqrt{\alpha^{\prime}}\ddot{\Phi}}{e^{\Phi\left(\tau\right)}\beta_{i}}\stackrel{{\scriptstyle[}}{{k}}=1]{\infty}{\sum}\frac{b_{k}}{1+\sigma^{2k-2}}$ (23)
$-\frac{\sqrt{2d}\dot{\Phi}}{e^{\Phi\left(\tau\right)}\beta_{i}}\stackrel{{\scriptstyle[}}{{k}}=1]{\infty}{\sum}\frac{k\sigma^{2k-1}b_{k+1}}{\left(1+\sigma^{2k}\right)^{2}},$
where $b_{k}\left(\beta_{i}\right)$ is a constant and satisfies the following constraint:
$\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}b_{1}\left(\beta_{i}\right)=2,\qquad\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}b_{k}\left(\beta_{i}\right)=0,\qquad k>1.$ (24)
The detailed expressions for $b_{k}\left(\beta_{i}\right)$ can be found in the Appendix. This solution is non-perturbative because it holds for any value of $\alpha^{\prime}$ and includes all-order higher curvature corrections. Furthermore, the associated Kretschmann scalar is given by
$R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ $=$ $2\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}^{2}\right)^{2}+2\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}^{4}\right)$ (25)
$+8\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}^{2}\dot{H}_{i}\right)+4\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}\dot{H}_{i}^{2}\right)$
$=$ $\left(\stackrel{{\scriptstyle[}}{{k}}=0]{N}{\sum}\lambda_{k}\sigma^{2k}\right)^{-\varrho}F\left(\tau\right),$
where $\varrho$ is some positive number and $F\left(\tau\right)$ is a regular function. So the singularities for the spacetime and dilaton appear if and only if
$\mathrm{Singular\;condition}:\qquad\stackrel{{\scriptstyle[}}{{k}}=0]{N}{\sum}\lambda_{k}\sigma^{2k}=0,$ (26)
has real roots. Specifically, to match the recently known coefficients $c_{4,0}$ and $c_{4,1}$ for the 4th order of $\alpha^{\prime}$ correction, and set $d=3$, we obtain
$\stackrel{{\scriptstyle[}}{{k}}=0]{N}{\sum}\lambda_{k}\sigma^{2k}$ $=$ $2099520\sigma^{6}+1283040\sigma^{4}+13680\sigma^{2}$ (27)
$+15336\zeta\left(3\right)+61187.$
This equation has no real roots for $\sigma$, indicating that the singularities of the Schwarzschild and rotating, accreting black holes are eliminated in the non-perturbative solution (see Appendix for details). Indeed, the parameter $\lambda_{k}$ is determined by the condition $c_{n\leq k}$. Considering any order $k$, although $\lambda_{n\leq k}$ are fixed by the coefficients $c_{n\leq k}$, we always have freedom to choose $\lambda_{n>k}$ which violate the singular condition (26), and obtaining the non-singular solutions. Moreover, It is straightforward to verify that this solution (23) covers the perturbative solution (LABEL:eq:perturbed_solution) as $\alpha^{\prime}\rightarrow 0$. Finally, it is worth noting that string effects only resolve the singularity but do not modify the physical properties of black holes beyond the near-singular region.
In conclusion, we demonstrated that the near-singularity geometry of Einstein’s black holes can be effectively described by the Kasner universe. Since, the Kasner universe only depends on $\tau$, we used the Hohm-Zwiebach action to study the singularity problem of black holes. The results indicated that the singularities of Schwarzschild and rotating, accreting black holes may be resolved by considering the $\alpha^{\prime}$ corrections of string theory if the singular condition (26) is broken by the known coefficients $c_{k}$. Nevertheless, we can still make the solution non-singular by adjusting $\lambda_{n>k}$ freely [15]. Moreover, according to the BKL proposal, the general realistic collapse to a spacelike singularity can be described by the Kasner universe, which consists of different Kasner epochs characterized by various choices of the Kasner exponents $\beta_{i}$ [29]. Since our results can apply to any $\beta_{i}$, it suggests that it is possible to study the singularity problem for the astrophysically realistic black holes through string theory.
Acknowledgements.
Acknowledgments.
We would especially like to thank Maurizio Gasperini, Robert Mann, Gabriele Veneziano for reading the draft and providing valuable suggestions and insights. We are also deeply indebt to Ronak Soni, Aron Wall, Peng Wang, Haitang Yang for many illuminating discussions and suggestions. HW is partially supported by NSFC Grant No.12275183, 12275184, 12105191 and 11875196. ZY is supported by an Internal Graduate Studentship of Trinity College, Cambridge and the AFOSR grant FA9550-19-1-0260 “Tensor Networks and Holographic Spacetime”. SY is supported by NSFC Grant No.12105031 and cstc2021jcyj-bshX0227.
Appendix
Here, we present the complete non-perturbative solution for the anisotropic HZ action. Considering the following ansatz:
$ds^{2}=-d\tau^{2}+\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}a_{i}\left(\tau\right)^{2}dx_{i}^{2},$ (28)
the EOM for the anisotropic HZ action can be given as:
$\ddot{\Phi}+\frac{1}{2}\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}f_{i}\left(H_{i}\right)$ $=$ $0,$
$\left(\frac{d}{d\tau}-\dot{\Phi}^{2}\right)f_{i}\left(H_{i}\right)$ $=$ $0,$
$\dot{\Phi}^{2}+\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}g_{i}\left(H_{i}\right)$ $=$ $0,$ (29)
.1 Perturbative solutions up to an arbitrary higher order $\alpha^{\prime n}$:
The corresponding perturbative solution can be calculated order by order, and it is given by:
$\Phi\left(\tau\right)$ $=$ $-\log\left(\gamma\tau\right)-\frac{\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}}{4}\frac{\alpha^{\prime}}{\tau^{2}}+\frac{9\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)^{2}+16\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{6}}{144}\frac{\alpha^{\prime 2}}{\tau^{4}}$
$-\frac{1}{720}\left[64\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)\left(\stackrel{{\scriptstyle[}}{{k}}=1]{d}{\sum}\beta_{k}^{6}\right)+15\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)^{3}+72\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{8}\right.$
$\left.+73728\left(\stackrel{{\scriptstyle[}}{{k}}=1]{d}{\sum}\beta_{k}^{4}\right)^{2}c_{41}+36864\left(\stackrel{{\scriptstyle[}}{{k}}=1]{d}{\sum}\beta_{k}^{8}\right)c_{40}\right]\frac{\alpha^{\prime 3}}{\tau^{6}}+\cdots,$
$H_{i}\left(\tau\right)$ $=$ $\frac{\beta_{i}}{\tau}-\frac{\beta_{i}\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}+4\beta_{i}^{2}\right)}{4\tau^{3}}\alpha^{\prime}+\frac{\text{$\beta_{i}$}\left(576\beta_{i}^{4}+216\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\beta_{i}^{2}+32\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{6}+27\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)^{2}\right)}{288\tau^{5}}\alpha^{\prime 2}$ (30)
$-\frac{\beta_{i}\alpha^{\prime 3}}{1920\tau^{7}}\left[98304\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{8}+40\beta_{i}^{6}\right)c_{4,0}+196608\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\left(\stackrel{{\scriptstyle[}}{{k}}=1]{d}{\sum}\beta_{k}^{4}+40\beta_{i}^{2}\right)c_{4,1}\right.$
$+7680\beta_{i}^{6}+4800\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)\beta_{i}^{4}+640\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{6}\right)\beta_{i}^{2}+900\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)^{2}\beta_{i}^{2}$
$\left.+224\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)\left(\stackrel{{\scriptstyle[}}{{k}}=1]{d}{\sum}\beta_{k}^{6}\right)+75\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)^{3}+192\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{8}\right]+\cdots,$
where $\gamma$ is an integration constant. Note that this solution can be easily generalized to an arbitrary higher order using Mathematica.
.2 Non-perturbative solutions which cover $\alpha^{\prime n}$ as $\alpha^{\prime}\rightarrow 0$:
The non-perturbative solution of EOM (29) is given by:
$f_{i}\left(H_{i}\right)$ $=$ $-\frac{2\beta_{i}}{\sqrt{\alpha^{\prime}}}\left(\stackrel{{\scriptstyle[}}{{k}}=0]{N}{\sum}\lambda_{k}\sigma^{2k}\right)^{-\frac{1}{2N}},$
$g_{i}\left(H_{i}\right)$ $=$ $-\frac{d}{2\alpha^{\prime}N^{2}}\left(\frac{\stackrel{{\scriptstyle[}}{{k}}=1]{N}{\sum}2k\lambda_{k}\sigma^{2k-1}}{\stackrel{{\scriptstyle[}}{{k}}=0]{N}{\sum}\lambda_{k}\sigma^{2k}}\right)^{2}\stackrel{{\scriptstyle[}}{{k}}=1]{\infty}{\sum}\frac{b_{k}}{1+\sigma^{2k-2}},$
$H_{i}$ $=$ $\frac{2d}{2N\beta_{i}\sqrt{\alpha^{\prime}}\left(\stackrel{{\scriptstyle[}}{{k}}=0]{N}{\sum}\lambda_{k}\sigma^{2k}\right)^{2-\frac{1}{2N}}}\left[-\left(\stackrel{{\scriptstyle[}}{{k}}=1]{N}{\sum}2k\left(2k-1\right)\lambda_{k}\sigma^{2k-2}\right)\left(\stackrel{{\scriptstyle[}}{{k}}=0]{N}{\sum}\lambda_{k}\sigma^{2k}\right)\stackrel{{\scriptstyle[}}{{k}}=1]{\infty}{\sum}\frac{b_{k}}{1+\sigma^{2k-2}}\right.$ (31)
$\left.+\left(\stackrel{{\scriptstyle[}}{{k}}=1]{N}{\sum}2k\lambda_{k}\sigma^{2k-1}\right)^{2}\stackrel{{\scriptstyle[}}{{k}}=1]{\infty}{\sum}\frac{b_{k}}{1+\sigma^{2k-2}}+\left(\stackrel{{\scriptstyle[}}{{k}}=1]{N}{\sum}2k\lambda_{k}\sigma^{2k-1}\right)\stackrel{{\scriptstyle[}}{{k}}=0]{N}{\sum}\lambda_{k}\sigma^{2k}\stackrel{{\scriptstyle[}}{{k}}=1]{\infty}{\sum}\frac{kb_{k+1}\sigma^{2k-1}}{\left(1+\sigma^{2k}\right)^{2}}\right].$
To compare this result with (30) as $\alpha^{\prime}\rightarrow 0$, we can fix $\gamma=\frac{1}{\sqrt{\alpha^{\prime}}}$ and the coefficients $\lambda_{k}$ and $b_{k}$:
$\lambda_{N}$ $=$ $\left(\frac{1}{2d}\right)^{N},$
$\lambda_{N-1}$ $=$ $dN\left(\frac{1}{2d}\right)^{N}\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4},$
$\lambda_{N-2}$ $=$ $\left(\frac{1}{2}\left(N-1\right)\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)^{2}-\frac{8}{9}\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{6}\right)d^{2}N\left(\frac{1}{2d}\right)^{N},$
$\lambda_{N-3}$ $=$ $\left[\left(\frac{64}{45}-\frac{8}{9}N\right)\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{6}\stackrel{{\scriptstyle[}}{{k}}=1]{d}{\sum}\beta_{k}^{4}+\frac{8}{5}\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{8}+\frac{N^{2}-3N+2}{6}\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)^{3}\right.$
$\left.+\frac{8192}{5}\left(\stackrel{{\scriptstyle[}}{{k}}=1]{d}{\sum}\beta_{k}^{4}\right)^{2}c_{4,1}+\frac{4096}{5}\left(\stackrel{{\scriptstyle[}}{{k}}=1]{d}{\sum}\beta_{k}^{8}\right)c_{4,0}\right]d^{3}N\left(\frac{1}{2d}\right)^{N},$
$\ldots$ (32)
and
$b_{1}\left(\beta_{i}\right)$ $=$ $2\beta_{i}^{2},$
$b_{2}\left(\beta_{i}\right)$ $=$ $-d\beta_{i}^{2}\left(\beta_{i}^{2}-\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right),$
$b_{3}\left(\beta_{i}\right)$ $=$ $\frac{d\beta_{i}^{2}}{3}\left(3\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{k}^{4}-8d\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{6}+8d\beta_{i}^{4}-3\beta_{i}^{2}\right),$
$b_{4}\left(\beta_{i}\right)$ $=$ $\frac{d\beta_{i}^{2}}{9}\left[8d^{2}\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{6}\right)\left(\stackrel{{\scriptstyle[}}{{k}}=1]{d}{\sum}\beta_{k}^{4}\right)+72d^{2}\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{8}\right)-9\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)\right.$
$+\beta_{i}^{2}\left(16d^{2}\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{6}\right)+9\right)-24d^{2}\beta_{i}^{4}\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)-72d^{2}\beta_{i}^{6}$
$\left.+73728d^{2}c_{4,1}\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)\left(\left(\stackrel{{\scriptstyle[}}{{k}}=1]{d}{\sum}\beta_{k}^{4}\right)-\beta_{i}^{2}\right)+36864d^{2}c_{4,0}\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{8}-\beta_{i}^{6}\right)\right],$
$\ldots$,… , (33)
where
$\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}b_{1}\left(\beta_{i}\right)=2,\qquad\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}b_{k}\left(\beta_{i}\right)=0,\qquad k>1,$ (34)
Now, we can study the singularity of the non-perturbative solution (31) utilizing the associated Kretschmann scalar:
$R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ $=$ $2\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}^{2}\right)^{2}+2\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}^{4}\right)+8\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}^{2}\dot{H}_{i}\right)+4\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}\dot{H}_{i}^{2}\right)$ (35)
$=$ $\left(\stackrel{{\scriptstyle[}}{{k}}=0]{N}{\sum}\lambda_{k}\sigma^{2k}\right)^{-\varrho}F\left(\tau\right),$
where $\varrho$ is some positive number and $F\left(\tau\right)$ is a regular function. So, the singularities for the spacetime and dilaton appear if and only if:
$\stackrel{{\scriptstyle[}}{{k}}=0]{N}{\sum}\lambda_{k}\sigma^{2k}=0,$ (36)
has real roots.
.3 Non-perturbative solutions which cover $\alpha^{\prime 3}$ as $\alpha^{\prime}\rightarrow 0$ (special example):
To cover the perturbative solution (30) up to the order of $\alpha^{\prime 3}$ as $\alpha^{\prime}\rightarrow 0$, and including the known coefficients $c_{1,0}=-\frac{1}{8}$, $c_{2,0}=\frac{1}{64}$, $c_{3,0}=-\frac{1}{3.2^{7}}$, $c_{4,0}=\frac{1}{2^{12}}-\frac{3}{2^{12}}\zeta\left(3\right)$, $c_{4,1}=\frac{1}{2^{16}}+\frac{1}{2^{12}}\zeta\left(3\right)$, we need to choose $N=3$ and $d=3$, which denotes the singularity regions of the four-dimensional Schwarzschild and rotating, accreting black holes. The corresponding non-perturbative solution (31) is given by:
$\Phi\left(\tau\right)$ $=$ $-\frac{1}{6}\log\left(\lambda_{0}+\lambda_{1}\sigma^{2}+\lambda_{2}\sigma^{4}+\lambda_{3}\sigma^{6}\right),\qquad\sigma^{2}\equiv 6\frac{\tau^{2}}{\alpha^{\prime}},$
$f_{i}\left(H_{i}\right)$ $=$ $-\frac{2\beta_{i}}{\sqrt{\alpha^{\prime}}}\left(\lambda_{0}+\lambda_{1}\sigma^{2}+\lambda_{2}\sigma^{4}+\lambda_{3}\sigma^{6}\right)^{-\frac{1}{6}},$
$g_{i}\left(H_{i}\right)$ $=$ $-\frac{\left(\lambda_{1}\sigma+2\lambda_{2}\sigma^{3}+3\lambda_{3}\sigma^{5}\right)^{2}\left(\frac{b_{1}}{2}+\frac{b_{2}}{1+\sigma^{2}}+\frac{b_{3}}{1+\sigma^{4}}+\frac{b_{4}}{1+\sigma^{6}}\right)}{3\alpha^{\prime}\left(\lambda_{0}+\lambda_{1}\sigma^{2}+\lambda_{2}\sigma^{4}+\lambda_{3}\sigma^{6}\right)^{2}},$
$H_{i}\left(\tau\right)$ $=$ $\frac{1}{\sqrt{\alpha^{\prime}}\beta_{i}}\left(\lambda_{0}+\lambda_{1}\sigma^{2}+\lambda_{2}\sigma^{4}+\lambda_{3}\sigma^{6}\right)^{-\frac{11}{6}}\left[-\left(\lambda_{0}+\lambda_{1}\sigma^{2}+\lambda_{2}\sigma^{4}+\lambda_{3}\sigma^{6}\right)\times\right.$ (37)
$\left(2\lambda_{1}+12\lambda_{2}\sigma^{2}+30\lambda_{3}\sigma^{3}\right)\left(\frac{b_{1}}{2}+\frac{b_{2}}{1+\sigma^{2}}+\frac{b_{3}}{1+\sigma^{4}}+\frac{b_{4}}{1+\sigma^{6}}\right)$
$+\left(2\lambda_{1}\sigma+4\lambda_{2}\sigma^{3}+6\lambda_{3}\sigma^{5}\right)\left(\lambda_{0}+\lambda_{1}\sigma^{2}+\lambda_{2}\sigma^{4}+\lambda_{3}\sigma^{6}\right)\times$
$\left(\frac{\sigma b_{2}}{\left(1+\sigma^{2}\right)^{2}}+\frac{2\sigma^{3}b_{3}}{\left(1+\sigma^{4}\right)^{2}}+\frac{3\sigma^{5}b_{4}}{\left(1+\sigma^{6}\right)^{2}}\right)+\left(2\lambda_{1}\sigma+4\lambda_{2}\sigma^{3}+6\lambda_{3}\sigma^{5}\right)^{2}\times$
$\left.\left(\frac{b_{1}}{2}+\frac{b_{2}}{1+\sigma^{2}}+\frac{b_{3}}{1+\sigma^{4}}+\frac{b_{4}}{1+\sigma^{6}}\right)\right],$
[…]
$b_{1}\left(\beta_{i}\right)$ $=$ $2\beta_{i}^{2},$
$b_{2}\left(\beta_{i}\right)$ $=$ $3\beta_{i}^{2}\left(\frac{11}{27}-\beta_{i}^{2}\right),$
$b_{3}\left(\beta_{i}\right)$ $=$ $\beta_{i}^{2}\left(24\beta_{i}^{4}-3\beta_{i}^{2}-\frac{245}{81}\right),$
$b_{4}\left(\beta_{i}\right)$ $=$ $\frac{\beta_{i}^{2}}{3}\left(270336c_{41}\left(\frac{11}{27}-\beta_{i}^{2}\right)+331776c_{40}\left(\frac{19}{243}-\beta_{i}^{6}\right)-648\beta_{i}^{6}-88\beta_{i}^{4}+\frac{931\beta_{i}^{2}}{27}+\frac{38047}{729}\right),$ (38)
where $c_{4,0}=\frac{1}{2^{12}}-\frac{3}{2^{12}}\zeta\left(3\right)$ and $c_{4,1}=\frac{1}{2^{16}}+\frac{1}{2^{12}}\zeta\left(3\right)$. It is worth noting that
$\stackrel{{\scriptstyle[}}{{i}}=1]{3}{\sum}b_{1}\left(\beta_{i}\right)=2,$ (39)
and
$\stackrel{{\scriptstyle[}}{{i}}=1]{3}{\sum}b_{2}\left(\beta_{i}\right)=\stackrel{{\scriptstyle[}}{{i}}=1]{3}{\sum}b_{3}\left(\beta_{i}\right)=\stackrel{{\scriptstyle[}}{{i}}=1]{3}{\sum}b_{4}\left(\beta_{i}\right)=0.$ (40)
The corresponding Kretschmann scalar is:
$R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ $=$ $2\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}^{2}\right)^{2}+2\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}^{4}\right)+8\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}^{2}\dot{H}_{i}\right)+4\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}\dot{H}_{i}^{2}\right)$ (41)
$=$ $\left(\left(15336\zeta\left(3\right)+61187\right)+13680\sigma^{2}+1283040\sigma^{4}+2099520\sigma^{6}\right)^{-\varrho}F\left(\tau\right),$
Then, it is easy to check that this equation has no real root, implying the singularities of four-dimensional Schwarzschild and rotating, accreting black holes are removed. Finally, we wish to note that although the coefficients $c_{k\geq 5}$ are unknown, in principle, they can be determined perturbatively by the worldsheet anomaly cancellation.