### v3

# Revisiting Schwarzschild black hole singularity through string theory

Abstract
In this letter, we derive the singular condition for black holes and demonstrate the potential resolution of the Schwarzschild black hole singularity in general relativity using non-perturbative $\alpha^{\prime}$ corrections of string theory. This work is motivated by the Belinskii, Khalatnikov and Lifshitz (BKL) proposal, which suggests that the structure of the black hole interior in vacuum Einstein’s equations can be transformed into the Kasner universe near the singularity. This transformation allows for the description of the black hole interior using the $O\left(d,d\right)$ invariant anisotropic Hohm-Zwiebach action, which includes all orders of $\alpha^{\prime}$ corrections.
preprint: APS/123-QED
Black holes are the most important objects in our universe, as understanding their internal structure and microstate could reveal the fundamental nature of physics. However, gravitational collapse inevitably leads to a curvature singularity at the center of a black hole, a feature that remains unavoidable in the framework of Einstein’s gravity [1, 2]. This suggests that general relativity is incomplete and breaks down in the vicinity of this singularity. String theory, a candidate theory of quantum gravity, is expected to modify the general relativity by including higher-derivative corrections and resolving this longstanding issue. However, to date, it has only provided insights into two-dimensional string black holes [3, 4], leaving unresolved the singularity problem of the four-dimensional black holes of Einstein’s gravity. This unbridgeable gap arises from the unknown non-perturbative aspects of string theory when approaching the curvature singularity.
In this letter, we aim to investigate whether the singularity of the Schwarzschild and rotating, accreting black holes can be resolved by higher-derivative $\alpha^{\prime}$ corrections of string theory. This work is motivated by the recent progress in classifying all orders of $\alpha^{\prime}$ corrections for specific backgrounds by Hohm and Zwiebach [5, 6, 7]. In previous work, Sen has demonstrated that when the classical fields, including the metric, Kalb-Ramond field and dilaton, are independent of $m$ coordinates, the field theory exhibits an $O\left(m,m\right)$ symmetry to all orders in $\alpha^{\prime}$ [8, 9]. This result has been verified for a few orders in $\alpha^{\prime}$, indicating that the standard form of $O\left(d,d\right)$ transformations can be preserved through suitable field redefinitions, even in the presence of $\alpha^{\prime}$ corrections [10, 8, 9, 11, 12, 13]. Based on these results, Hohm and Zwiebach conjectured that the standard form of $O\left(d,d\right)$ transformations also preserves for all orders in $\alpha^{\prime}$, enabling the classification of all $O\left(d,d\right)$ invariant $\alpha^{\prime}$ corrections. Since the corresponding $O\left(d,d\right)$ invariant action involves only first-order derivatives, it can be exactly solved. Consequently, it raises the question of whether black hole singularities can be cured based on this new development. Recent studies have successfully resolved the big bang singularity [14, 15, 16, 17, 18, 19] and the curvature singularities of two-dimensional string black holes [20, 21, 22, 23] in this framework. However, this approach still cannot resolve the singularity of the Schwarzschild and rotating, accreting black holes. The challenge arises from the fact that the spherical part of the black hole metric does not exhibit $O\left(d,d\right)$ symmetry. Therefore, directly applying the Hohm-Zwiebach action becomes impossible.
Encouragingly, the general structure near the black hole singularity, namely Kasner universe, exhibits $O\left(d,d\right)$ symmetry, opening avenues for understanding the black hole singularity in non-perturbative string theory. To be specific, let us recall the tree-level string effective action:
$I_{\mathrm{String}}=\frac{1}{16\pi G_{D}}\int d^{D}x\sqrt{-g}e^{-2\phi}\left(R+4\left(\partial\phi\right)^{2}\right).$ (1)
where we adopt $16\pi G_{D}=1$, $\phi\left(x\right)$ denotes a physical dilaton and we set Kalb-Ramond field $b_{\mu\nu}=0$ for simplicity. Moreover, when $\phi$ is a constant, the action reduces to the Einstein-Hilbert action and we will consider the stringy corrections to this action in this letter. On the other hand, since the spacetime dimension of bosonic string theory is fixed to be $26$, we can see the manifold as a direct product of the black hole and a compact internal space [24]. Due to the fact that the decoupling of the black hole and internal space hold for all order of $\alpha^{\prime}$ corrections [25], we only need to consider the black hole part in the following calculations. To obtain the Kasner universe near the black hole singularity, we first consider the $D=n+3$ dimensional Schwarzschild-Tangherlini black hole [26, 27]:
$ds^{2}=-\left(1-\left(\frac{r_{0}}{r}\right)^{n}\right)dt^{2}+\frac{dr^{2}}{1-\left(\frac{r_{0}}{r}\right)^{n}}+r^{2}d\Omega_{n+1}^{2},$ (2)
where $d\Omega_{n+1}^{2}$ denotes the metric of an $\left(n+1\right)$ dimensional sphere, and $r_{0}$ is an event horizon. Inside the event horizon, the metric can be transformed into the Kasner metric near the singularity [28]:
$ds^{2}=-d\tau^{2}+\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\tau^{2\beta_{i}}dx_{i}^{2}.$ (3)
The spacelike singularity is now located at $\tau=0$ which corresponds to $r=0$ of Schwarzschild black hole. Moreover, in order to satisfy the vacuum Einstein’s equations, the Kasner exponents $\beta_{i}$ must satisfy the following constraints
$\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\beta_{i}^{2}=1,\qquad\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\beta_{i}=1.$ (4)
For the four-dimensional Schwarzschild black hole, we have $\left(\beta_{1},\beta_{2},\beta_{3}\right)=\left(-\frac{1}{3},\frac{2}{3},\frac{2}{3}\right)$. One may wonder whether the initial spherical symmetry is broken in the Kasner metric (3). Near the singularity, the curvature becomes large enough that the evolution of the black hole geometry at different spatial points decouples. It means that the partial differential Einstein’s equations become ordinary differential equations (ODEs) with respect to time [29]. Therefore, there is no reason to assume the initial spherical symmetry is also maintained in ODEs’ solution. The BKL approximation (with the Kasner metric as its solution) is sufficient to describe this region. It depicts an asymmetrically spherical shell chaotically collapsing near the singularity.
Apart from the Schwarzschild black hole, the Kerr black hole is highly interesting. However, the presence of an inner horizon in a rotating black hole transforms the spacelike singularity into a timelike singularity. Poisson and Israel studied the mass inflation instability at the inner horizon, which leads to the formation of a null singularity at the inner horizon [30]. In subsequent work, Burko found that if the radiation power law does not drop off quickly, a spacelike singularity rather than a null singularity will be formed. In other words, the inner horizon can be superseded by the BKL singularity [31]. Hamilton then numerically studied the evolution of the inner horizon of a rotating, accreting black hole, which implies that the spacetime almost undergoes a BKL collapse [32]. Subsequently, the analytic model of this collapse, called the inflationary Kasner solution, was constructed, including two Kasner epochs: an inflation epoch with Kasner exponents $\left(\beta_{1},\beta_{2},\beta_{3}\right)=\left(1,0,0\right)$ and a collapse epoch characterized by Kasner exponents $\left(\beta_{1},\beta_{2},\beta_{3}\right)=\left(-\frac{1}{3},\frac{2}{3},\frac{2}{3}\right)$ approaching the spacelike singularity, which approximates a Schwarzschild geometry [33].
Since the Kasner universe only depends on $\tau$, the string action in this background possesses the $O\left(d,d\right)$ symmetry. This allows us to study the effects of string theory non-perturbatively in near singularity region and examine how it may potentially resolve the black hole singularities (the previous discussions on anisotropic Hohm-Zwiebach action can be found in references [34, 35, 17, 36]. Although the spacelike singularity of isotropic string cosmology has been resolved in many references [37, 38], the anisotropic case is highly non-trivial and lacks any progress on its exact solution [36]). Thus, let us use the following ansatz:
$ds^{2}$ $=$ $-d\tau^{2}+g_{ij}dx^{i}dx^{j},$
$g_{ij}$ $=$ $\mathrm{diag}\left(a_{1}\left(\tau\right)^{2},a_{2}\left(\tau\right)^{2},\ldots,a_{i}\left(\tau\right)^{2}\right),$ (5)
where $i=1,\ldots,d$. The corresponding anisotropic Hohm-Zwiebach action is given by:
$I_{\mathrm{HZ}}$ $=$ $\int d^{D}x\sqrt{-g}e^{-2\phi}\left(R+4\left(\partial\phi\right)^{2}\right$. (6)
$\left.+\frac{1}{4}\alpha^{\prime}\left(R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}+\cdots\right)+\alpha^{\prime 2}\left(\cdots\right)+\cdots\right)$
$=$ $\int d\tau e^{-\Phi}\left(-\dot{\Phi}^{2}-\frac{1}{8}\mathrm{Tr}\left(\dot{\mathcal{S}}^{2}\right)\right$.
$+\alpha^{\prime}c_{2,0}\mathrm{Tr}\left(\dot{\mathcal{S}}^{4}\right)+\alpha^{\prime 2}c_{3,0}\mathrm{Tr}\left(\dot{\mathcal{S}}^{6}\right)$
$+\alpha^{\prime 3}\left[c_{4,0}\mathrm{Tr}\left(\dot{\mathcal{S}}^{8}\right)+c_{4,1}\left(\mathrm{Tr}\left(\dot{\mathcal{S}}^{4}\right)\right)^{2}\right]$
[…]
where $\dot{f}\left(\tau\right)\equiv\partial_{\tau}f\left(\tau\right)$. Moreover, there is no arbitrariness for this action as a string effective action due to the form of this action is constrained by the symmetry and has removed ambiguous coefficients through the field redefinition. On the other hand, it has been verified that at order $\alpha^{\prime 3}$ of Type II string theory [39] and order $\alpha^{\prime}$ of general torus compactifications [40], they are consistent with (6) with specific coefficients $c_{i,j}$. The string vacuum is also a solution of this action [6, 16]. For a bosonic case, $c_{1,0}=-\frac{1}{8}$, $c_{2,0}=\frac{1}{64}$, $c_{3,0}=-\frac{1}{3.2^{7}}$, $c_{4,0}=\frac{1}{2^{12}}-\frac{3}{2^{12}}\zeta\left(3\right)$, $c_{4,1}=\frac{1}{2^{16}}+\frac{1}{2^{12}}\zeta\left(3\right)$ and $c_{k>4}$’s are unknown [13]. This action systematically provides all-order higher curvature corrections to Einstein’s gravity. For example, the leading order is the Einstein-Hilbert action ($\phi=\mathrm{constant}$), and the first order of $\alpha^{\prime}$ correction is the Gauss-Bonnet term. Moreover, $O\left(d,d\right)$ invariant dilaton is defined by $\Phi\equiv 2\phi-\log\sqrt{-g}$. The matrix $\mathcal{S}$ is defined as
$\mathcal{S}\equiv\left(\begin{array}[]{cc}0&g_{ij}\\ g_{ij}^{-1}&0\end{array}\right),\quad\mathrm{Tr}\left(\dot{\mathcal{S}}^{2k}\right)=\left(-4\right)^{k}2\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}^{2k},$ (7)
where $H_{i}\equiv\frac{\dot{a}_{i}\left(\tau\right)}{a_{i}\left(\tau\right)}$ denotes the Hubble parameter. In addition, the $O\left(d,d\right)$ symmetry of the action (6) is manifested by the following transformations:
$\Phi\rightarrow\Phi,\qquad a_{i}\rightarrow a_{i}^{-1},\qquad H_{i}\rightarrow-H_{i}.$ (8)
However, it is challenging to obtain the solutions of EOM from the action (6) due to the multi-trace terms. As an example, let us consider the action at the order of $\alpha^{\prime 3}$ which introduces the first multi-trace term into the action (6). The action can be written as follows:
$I_{\mathrm{HZ}}^{\left(3\right)}=\alpha^{\prime 3}\left[c_{4,0}\mathrm{Tr}\left(\dot{\mathcal{S}}^{8}\right)+c_{4,1}\left(\mathrm{Tr}\left(\dot{\mathcal{S}}^{4}\right)\right)^{2}\right].$ (9)
For the isotropic background:
$I_{\mathrm{HZ}}^{\left(3\right)}$ $=$ $\alpha^{\prime 3}\left(c_{4,0}\left(-4\right)^{4}2dH^{8}+c_{4,1}\left(-4\right)^{4}4d^{2}H^{8}\right)$ (10)
$=$ $\alpha^{\prime 3}\left(\mathrm{constant}\times H^{8}\right),$
all orders of $\alpha^{\prime}$ corrections can be straightforwardly summed in a simple manner proportional to $H^{2k}$. And it is easy to obtain the corresponding EOM. But, for the anisotropic case:
$I_{\mathrm{HZ}}^{\left(3\right)}=\alpha^{\prime 3}\left(-4\right)^{4}\left[2c_{4,0}\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}^{8}+4c_{4,1}\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}^{4}\right)^{2}\right],$ (11)
which cannot be summed together and the effects of multi-trace cannot be neglected.
The aim of this letter is to calculate the non-perturbative and non-singular black hole solution of the action (6) with all-order higher curvature corrections. To calculate its solution, we first recall the EOM of the action (6):
$\ddot{\Phi}+\frac{1}{2}\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}f_{i}\left(H_{i}\right)$ $=$ $0,$
$\left(\frac{d}{d\tau}-\dot{\Phi}\right)f_{i}\left(H_{i}\right)$ $=$ $0,$
$\dot{\Phi}^{2}+\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}g_{i}\left(H_{i}\right)$ $=$ $0,$ (12)
where
$f_{i}\left(H_{i}\right)$ $=$ $-2H_{i}-2\alpha^{\prime}H_{i}^{3}-2\alpha^{\prime 2}H_{i}^{5}$
$-8\alpha^{\prime 3}4^{4}\left[2c_{4,0}H_{i}^{7}+4c_{4,1}\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}H_{j}^{4}\right)H_{i}^{3}\right]+\cdots,$
$g_{i}\left(H_{i}\right)$ $=$ $-H_{i}^{2}-\frac{3}{2}\alpha^{\prime}H_{i}^{4}-\frac{5}{3}\alpha^{\prime 2}H_{i}^{6}$
$-7\alpha^{\prime 3}4^{4}\left[2c_{4,0}H_{i}^{8}+4c_{4,1}\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}H_{j}^{4}\right)H_{i}^{4}\right]+\cdots,$
and there is no additional constraint $\frac{d}{dH_{i}}g_{i}\left(H_{i}\right)=H_{i}\frac{d}{dH_{i}}f_{i}\left(H_{i}\right)$ for the anisotropic case. These EOM are consistent with refs. [35, 36]. When $\alpha^{\prime}=0$, the tree-level EOM (12) have a solution
$\Phi\left(\tau\right)$ $=$ $-\log\left(\tau-\tau_{0}\right)+\Phi_{0},$
$a_{i}\left(\tau\right)$ $=$ $C_{i}\left(\tau-\tau_{0}\right)^{\beta_{i}},$
$H_{i}\left(\tau\right)$ $=$ $\frac{\beta_{i}}{\tau-\tau_{0}},$ (14)
where $\tau_{0}$, $C_{i}$ and $\Phi_{0}$ are arbitrary constants. This result covers the solution of Einstein’s gravity (3) since the physical dilaton $\phi$ is a constant. The corresponding Kretschmann scalar is
$R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}=\frac{4}{\left(\tau-\tau_{0}\right)^{4}}\left(\frac{1}{2}\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\beta_{i}^{4}-2\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\beta_{i}^{3}+\frac{3}{2}\right).$ (15)
The spacelike singularity is located at $\tau=\tau_{0}$. Now, our aim is to remove this singularity using $\alpha^{\prime}$ corrections in string theory by the following strategy:
  • Firstly, we aim to calculate the perturbative solution of the EOM (12) in the perturbative regime as $\alpha^{\prime}\rightarrow 0$.
  • Then, in the non-perturbative regime for an arbitrary $\alpha^{\prime}$, we need to find a non-singular solution that satisfies the EOM (12) and covers arbitrary higher orders of perturbative solution with the multi-trace terms as $\alpha^{\prime}\rightarrow 0$ (the detailed calculations can be found in the Appendix).
To calculate the perturbation solution of the EOM (12), we introduce a new variable $\Omega$:
$\Omega\equiv e^{-\Phi},$ (16)
where $\dot{\Omega}=-\dot{\Phi}\Omega$ and $\ddot{\Omega}=\left(-\ddot{\Phi}+\dot{\Phi}^{2}\right)\Omega$. And the EOM (12) become
$\ddot{\Omega}-\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}h_{i}\left(H_{i}\right)\right)\Omega$ $=$ $0,$
$\frac{d}{d\tau}\left(\Omega f_{i}\left(H_{i}\right)\right)$ $=$ $0,$
$\dot{\Omega}^{2}+\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}g_{i}\left(H_{i}\right)\right)\Omega^{2}$ $=$ $0,$ (17)
where we define a new function
$h_{i}\left(H_{i}\right)\equiv\frac{1}{2}H_{i}f_{i}\left(H_{i}\right)-g\left(H_{i}\right)=\alpha^{\prime}\frac{1}{2}H_{i}^{4}+\cdots,$ (18)
Next, we assume the perturbative solutions of the EOM (17) to be:
$\Omega\left(\tau\right)$ $=$ $\Omega_{\left(0\right)}\left(\tau\right)+\alpha^{\prime}\Omega_{\left(1\right)}\left(\tau\right)+\alpha^{\prime 2}\Omega_{\left(2\right)}\left(\tau\right)+\cdots,$
$H_{i}\left(\tau\right)$ $=$ $H_{i\left(0\right)}\left(\tau\right)+\alpha^{\prime}H_{i\left(1\right)}\left(\tau\right)+\alpha^{\prime 2}H_{i\left(2\right)}\left(\tau\right)+\cdots,$
where we denote $\Omega_{i}$ and $H_{i}$ as the $i$th order of the perturbative solutions. Therefore, the perturbative solution can be calculated order by order, which is
$H_{i}\left(\tau\right)$ $=$ $\frac{\beta_{i}}{\tau-\tau_{0}}-\frac{4\beta_{i}^{3}+\beta_{i}\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}}{4\left(\tau-\tau_{0}\right)^{3}}\alpha^{\prime}+\cdots,$
$\Omega\left(\tau\right)$ $=$ $\gamma\left(\tau-\tau_{0}\right)+\frac{\gamma}{4\left(\tau-\tau_{0}\right)}\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\beta_{i}^{4}\right)\alpha^{\prime}+\cdots,$
where $\gamma$ is an integration constant. The leading-order of (LABEL:eq:perturbed_solution) covers the geometry near the Schwarzschild singularity (3). Moreover, due to equation (16), we also have:
$\Phi\left(\tau\right)=-\log\left(\gamma\left(\tau-\tau_{0}\right)\right)-\frac{1}{4\left(\tau-\tau_{0}\right)^{2}}\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\beta_{i}^{4}\right)\alpha^{\prime}+\cdots.$ (21)
Based on the perturbative solutions (LABEL:eq:perturbed_solution) and (21), we now construct a non-perturbative  [^1] Our statement that this solution is non-perturbative does not rely on the convergence of the series. Rather, it emphasizes that the logarithmic resummation captures contributions that cannot be represented by any finite-order polynomial in $\alpha^{\prime}$. In this sense, the solution extends into the non-perturbative regime, effectively incorporating all orders of $\alpha^{\prime}$ corrections. and non-singular solution that satisfies the EOM (12). Assuming the coefficients $c_{k\leq n}$ are known, the dilaton solution takes the form
$\Phi\left(\tau\right)=-\frac{1}{2N}\log\left(\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}\right),\qquad\sigma^{2}\equiv 2d\frac{\tau^{2}}{\alpha^{\prime}},$ (22)
where $N\geq n-1$ is an arbitrary integer corresponding to the truncation order of the $\alpha^{\prime}$ expansion, and $\lambda_{0}>0$, $\lambda_{n}>0$. The coefficients $\lambda_{N}$, $\lambda_{N-1}$, … , $\lambda_{N-n+1}$ are determined by $c_{1}$, … , $c_{n}$, while the remaining parameters $\lambda_{0}$, $\lambda_{1}$, … , $\lambda_{N-n}$ are free and can be chosen to violate the singularity condition (28), which we will introduce shortly. For example, consider the first two terms of the expansion of (22) in the perturbative regime $\tau\rightarrow\infty$ (or equivalently $\alpha^{\prime}\rightarrow 0$):
$\Phi\left(\frac{\tau}{\alpha^{\prime}}\rightarrow\infty\right)$ $=$ $-\frac{1}{2}\log\left(2d\frac{\tau^{2}}{\alpha^{\prime}}\lambda_{N}^{1/N}\right)$ (23)
$-\frac{1}{2N}\frac{\lambda_{N-1}}{\lambda_{N}}\frac{1}{2d}\frac{\alpha^{\prime}}{\tau^{2}}+\cdots.$
As discussed earlier, the $N$-th order in the $\alpha^{\prime}$ expansion becomes dominant, and the coefficients $\lambda_{N}$ and $\lambda_{N-1}$ can be fixed by matching to the perturbative coefficients $c_{1}$, $c_{2}$. Comparing this expansion with the perturbative solution (21), we determine
$\lambda_{N}$ $=$ $\gamma^{2N}\left(\frac{\alpha^{\prime}}{2d}\right)^{N},$
$\lambda_{N-1}$ $=$ $dN\gamma^{2N}\left(\frac{\alpha^{\prime}}{2d}\right)^{N}\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4},$ (24)
where we have set $\tau_{0}=0$ for simplicity, and used $c_{1}=-1/8$, $c_{2}=1/64$. Higher-order coefficients are listed in the Appendix B. The corresponding functions $f_{i}$, $g_{i}$ and $H_{i}$ are given by
$f_{i}$ $=$ $-\frac{2\beta_{i}}{\sqrt{\alpha^{\prime}}}e^{\Phi\left(\tau\right)},$
$g_{i}$ $=$ $-\dot{\Phi}^{2}\mathrel{{\mathop{\sum}\limits_{k=1}^{\infty}}}\frac{b_{k}}{1+\sigma^{2k-2}},$
$H_{i}$ $=$ $\frac{\sqrt{\alpha^{\prime}}\ddot{\Phi}}{e^{\Phi\left(\tau\right)}\beta_{i}}\mathrel{{\mathop{\sum}\limits_{k=1}^{\infty}}}\frac{b_{k}}{1+\sigma^{2k-2}}$ (25)
$-\frac{\sqrt{2d}\dot{\Phi}}{e^{\Phi\left(\tau\right)}\beta_{i}}\mathrel{{\mathop{\sum}\limits_{k=1}^{\infty}}}\frac{k\sigma^{2k-1}b_{k+1}}{\left(1+\sigma^{2k}\right)^{2}},$
where $b_{k}\left(\beta_{i}\right)$ is a constant and satisfies the following constraint:
$\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}b_{1}\left(\beta_{i}\right)=2,\qquad\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}b_{k}\left(\beta_{i}\right)=0,\qquad k>1.$ (26)
The detailed expressions for $b_{k}\left(\beta_{i}\right)$ can be found in the Appendix. This solution is non-perturbative because it holds for any value of $\alpha^{\prime}$ and includes all-order higher curvature corrections. Furthermore, the associated Kretschmann scalar is given by
$R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ $=$ $2\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}^{2}\right)^{2}+2\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}^{4}\right)$ (27)
$+8\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}^{2}\dot{H}_{i}\right)+4\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\dot{H}_{i}^{2}\right)$
$=$ $\left(\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}\right)^{-\varrho}F\left(\tau\right),$
where $\varrho\equiv 8-\frac{2}{N},$ and $F\left(\tau\right)$ denotes a regular function. The explicit form of $F\left(\tau\right)$ together with a detailed discussion of its possible singular behavior are given in Eq. (43) in Appendix B. So the singularities for the spacetime and dilaton appear if and only if
$\mathrm{Singular\;condition}:\qquad\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}=0,$ (28)
has real roots. Specifically, to match the recently known coefficients $c_{4,0}$ and $c_{4,1}$ for the 4th order of $\alpha^{\prime}$ correction, and set $d=3$, we obtain
$\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}$ $=$ $2099520\sigma^{6}+1283040\sigma^{4}+13680\sigma^{2}$ (29)
$+15336\zeta\left(3\right)+61187.$
This equation has no real roots for $\sigma$, indicating that the singularities of the Schwarzschild and rotating, accreting black holes are eliminated in the non-perturbative solution (see Appendix for details). Moreover, in this condition, both the string coupling $g_{s}=\exp\left(\phi\right)$ and the curvature invariant $R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ in the Einstein frame remain regular. The detailed analysis is provided in Appendix D. Indeed, since our ansatz (22) is constructed under the assumption that the coefficients $c_{k\leq n}$ are known, the parameter $\lambda_{k}$ is determined by the condition $c_{n\leq k}$. Considering any order $k$, although $\lambda_{n\leq k}$ are fixed by the coefficients $c_{n\leq k}$, we always have freedom to choose $\lambda_{n>k}$ which violate the singular condition (28), and obtaining the non-singular solutions. It is worth noting that this ansatz does not represent the unique or complete solution but rather provides a constructive framework to investigate how higher-order $\alpha^{\prime}$ corrections in the $O\left(d,d\right)$-invariant effective action may contribute to singularity resolution. Once the full set of coefficients $c_{k}$ becomes available from string theory (e.g., in heterotic or bosonic models), the solution would need to be correspondingly modified. Moreover, It is straightforward to verify that this solution (25) covers the perturbative solution (LABEL:eq:perturbed_solution) as $\alpha^{\prime}\rightarrow 0$. Finally, it is worth noting that our results focus on the resolution of the singularity. The key point is that the $\alpha^{\prime}$ corrections become significant only when the curvature grows sufficiently large. In the vicinity of the singularity, the inverse gradients of the background fields become comparable to the fundamental string length, such that $\alpha^{\prime}R\geq 1$, $\alpha^{\prime}\left(\nabla\phi\right)^{2}\geq 1$, and so on. In this high-curvature regime, the contributions from all orders of $\alpha^{\prime}$ corrections must be included. Our non-perturbative and non-singular solution is precisely obtained in this region. In the low-curvature regime, where $\alpha^{\prime}\rightarrow 0$, our non-perturbative solution can be expanded perturbatively and reproduces the familiar Schwarzschild black hole of Einstein’s gravity at leading order. Furthermore, we would like to emphasize that there are several indications that the $O\left(d,d\right)$ invariance may also be preserved in the low-curvature black-hole region, although this point is often overlooked in the literature. The Schwarzschild black hole (with vanishing dilaton) is T-dual to the string black hole (with a nontrivial dilaton), as observed in [41, 42, 43]. This duality has recently been successfully incorporated into the $O\left(d,d\right)$-invariant framework at large $D$ in our previous work [44]. It is therefore possible to investigate whether the $O\left(d,d\right)$ invariance can be preserved throughout the entire spacetime and whether the transition between the near-singularity (high-curvature) region and the asymptotic (low-curvature) black-hole region is smooth.
In conclusion, it is believed that the singularity geometry of the Schwarzschild black hole solution in Einstein’s gravity can be effectively described by the Kasner universe. Since the Schwarzschild black hole is a solution of the string low-energy effective action with a constant dilaton, this property allows us to consistently introduce higher-order $\alpha^{\prime}$ corrections, which modify the near-singularity geometry while preserving the Schwarzschild solution at leading order in the low-curvature regime. More specifically, since the Kasner universe only depends on $\tau$, we used the Hohm-Zwiebach action to study the singularity problem of black holes. The results indicated that the singularities of Schwarzschild and rotating, accreting black holes may be resolved by considering the $\alpha^{\prime}$ corrections of string theory if the singular condition (28) is broken by the known coefficients $c_{k}$. Nevertheless, we can still make the solution non-singular by adjusting $\lambda_{n>k}$ freely [15]. Moreover, according to the BKL proposal, the general realistic collapse to a spacelike singularity can be described by the Kasner universe, which consists of different Kasner epochs characterized by various choices of the Kasner exponents $\beta_{i}$ [29]. Since our results can apply to any $\beta_{i}$, it suggests that it is possible to study the singularity problem for the astrophysically realistic black holes through string theory.
Acknowledgements.
Acknowledgments.
We would especially like to thank Maurizio Gasperini, Robert Mann, Gabriele Veneziano for reading the draft and providing valuable suggestions and insights. We are also deeply indebt to Ronak Soni, Aron Wall, Peng Wang, Haitang Yang for many illuminating discussions and suggestions. HW is partially supported by NSFC Grant No.12275183, 12275184, 12105191 and 11875196. ZY is supported by an Internal Graduate Studentship of Trinity College, Cambridge and the AFOSR grant FA9550-19-1-0260 “Tensor Networks and Holographic Spacetime”. SY is supported by NSFC Grant No.12105031 and cstc2021jcyj-bshX0227.
Appendix
Here, we present the complete non-perturbative solution for the anisotropic HZ action. Considering the following ansatz:
$ds^{2}=-d\tau^{2}+\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}a_{i}\left(\tau\right)^{2}dx_{i}^{2},$ (30)
the EOM for the anisotropic HZ action can be given as:
$\ddot{\Phi}+\frac{1}{2}\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}f_{i}\left(H_{i}\right)$ $=$ $0,$
$\left(\frac{d}{d\tau}-\dot{\Phi}^{2}\right)f_{i}\left(H_{i}\right)$ $=$ $0,$
$\dot{\Phi}^{2}+\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}g_{i}\left(H_{i}\right)$ $=$ $0,$ (31)
.1 Perturbative solutions up to an arbitrary higher order $\alpha^{\prime n}$:
The corresponding perturbative solution can be calculated order by order, and it is given by:
$\Phi\left(\tau\right)$ $=$ $-\log\left(\gamma\tau\right)-\frac{\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}}{4}\frac{\alpha^{\prime}}{\tau^{2}}+\frac{9\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)^{2}+16\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{6}}{144}\frac{\alpha^{\prime 2}}{\tau^{4}}$
$-\frac{1}{720}\left[64\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)\left(\mathrel{{\mathop{\sum}\limits_{k=1}^{d}}}\beta_{k}^{6}\right)+15\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)^{3}+72\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{8}\right$.
$\left.+73728\left(\mathrel{{\mathop{\sum}\limits_{k=1}^{d}}}\beta_{k}^{4}\right)^{2}c_{41}+36864\left(\mathrel{{\mathop{\sum}\limits_{k=1}^{d}}}\beta_{k}^{8}\right)c_{40}\right]\frac{\alpha^{\prime 3}}{\tau^{6}}+\cdots,$
$H_{i}\left(\tau\right)$ $=$ $\frac{\beta_{i}}{\tau}-\frac{\beta_{i}\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}+4\beta_{i}^{2}\right)}{4\tau^{3}}\alpha^{\prime}+\frac{\text{$\beta_{i}$}\left(576\beta_{i}^{4}+216\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\beta_{i}^{2}+32\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{6}+27\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)^{2}\right)}{288\tau^{5}}\alpha^{\prime 2}$ (32)
$-\frac{\beta_{i}\alpha^{\prime 3}}{1920\tau^{7}}\left[98304\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{8}+40\beta_{i}^{6}\right)c_{4,0}+196608\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\left(\mathrel{{\mathop{\sum}\limits_{k=1}^{d}}}\beta_{k}^{4}+40\beta_{i}^{2}\right)c_{4,1}\right$.
$+7680\beta_{i}^{6}+4800\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)\beta_{i}^{4}+640\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{6}\right)\beta_{i}^{2}+900\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)^{2}\beta_{i}^{2}$
$\left.+224\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)\left(\mathrel{{\mathop{\sum}\limits_{k=1}^{d}}}\beta_{k}^{6}\right)+75\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)^{3}+192\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{8}\right]+\cdots,$
where $\gamma$ is an integration constant. Note that this solution can be easily generalized to an arbitrary higher order using Mathematica.
.2 Non-perturbative solutions which cover $\alpha^{\prime n}$ as $\alpha^{\prime}\rightarrow 0$:
The non-perturbative solution of EOM (31) is given by:
$\Phi\left(\tau\right)$ $=$ $-\frac{1}{2N}\log\left(\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}\right),\qquad\sigma^{2}\equiv 2d\frac{\tau^{2}}{\alpha^{\prime}},$
$f_{i}\left(H_{i}\right)$ $=$ $-\frac{2\beta_{i}}{\sqrt{\alpha^{\prime}}}\left(\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}\right)^{-\frac{1}{2N}},$
$g_{i}\left(H_{i}\right)$ $=$ $-\frac{d}{2\alpha^{\prime}N^{2}}\left(\frac{\mathrel{{\mathop{\sum}\limits_{k=1}^{N}}}2k\lambda_{k}\sigma^{2k-1}}{\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}}\right)^{2}\mathrel{{\mathop{\sum}\limits_{k=1}^{\infty}}}\frac{b_{k}}{1+\sigma^{2k-2}},$
$H_{i}$ $=$ $\frac{2d}{2N\beta_{i}\sqrt{\alpha^{\prime}}\left(\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}\right)^{2-\frac{1}{2N}}}\left[-\left(\mathrel{{\mathop{\sum}\limits_{k=1}^{N}}}2k\left(2k-1\right)\lambda_{k}\sigma^{2k-2}\right)\left(\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}\right)\mathrel{{\mathop{\sum}\limits_{k=1}^{\infty}}}\frac{b_{k}}{1+\sigma^{2k-2}}\right$. (33)
$\left.+\left(\mathrel{{\mathop{\sum}\limits_{k=1}^{N}}}2k\lambda_{k}\sigma^{2k-1}\right)^{2}\mathrel{{\mathop{\sum}\limits_{k=1}^{\infty}}}\frac{b_{k}}{1+\sigma^{2k-2}}+\left(\mathrel{{\mathop{\sum}\limits_{k=1}^{N}}}2k\lambda_{k}\sigma^{2k-1}\right)\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}\mathrel{{\mathop{\sum}\limits_{k=1}^{\infty}}}\frac{kb_{k+1}\sigma^{2k-1}}{\left(1+\sigma^{2k}\right)^{2}}\right].$
In the perturbative region, corresponding to $\tau\rightarrow\infty$ (or equivalently $\alpha^{\prime}\rightarrow 0$), the non-perturbative solution can be expanded as follows:
$\Phi\left(\frac{\tau}{\alpha^{\prime}}\rightarrow\infty\right)$ (34)
$=$ $-\frac{1}{2}\log\left(2d\frac{\tau^{2}}{\alpha^{\prime}}\lambda_{N}^{1/N}\right)-\frac{1}{2N}$
$\times\left(\frac{\lambda_{N-1}}{\lambda_{N}}\frac{1}{2d}\frac{\alpha^{\prime}}{\tau^{2}}+\frac{2\lambda_{N}\lambda_{N-2}-\lambda_{N-1}^{2}}{2\lambda_{N}^{2}}\frac{1}{\left(2d\right)^{2}}\frac{\alpha^{\prime 2}}{\tau^{4}}+\frac{3\lambda_{N}^{2}\lambda_{N-3}-3\lambda_{N}\lambda_{N-1}\lambda_{N-2}+\lambda_{N-1}^{3}}{3\lambda_{N}^{3}}\frac{1}{\left(2d\right)^{3}}\frac{\alpha^{\prime 3}}{\tau^{6}}\right$.
$\left.+\frac{4\lambda_{N}^{3}\lambda_{N-4}-2\lambda_{N}^{2}\lambda_{N-2}^{2}-4\lambda_{N}^{2}\lambda_{N-1}\lambda_{N-3}+4\lambda_{N}\lambda_{N-1}^{2}\lambda_{N-2}-\lambda_{N-1}^{4}}{4\lambda_{N}^{4}}\frac{1}{\left(2d\right)^{4}}\frac{\alpha^{\prime 4}}{\tau^{8}}+\mathcal{O}\left(\frac{1}{\tau^{10}}\right)\right),$
and
$H_{i}\left(\frac{\tau}{\alpha^{\prime}}\rightarrow\infty\right)$ (35)
$=$ $\frac{b_{1}d^{\frac{1}{2}}\lambda_{N}^{\frac{1}{2N}}}{\sqrt{2}\beta_{i}}\frac{1}{\tau}+\frac{\left(8N\lambda_{N}b_{2}-5\lambda_{N-1}b_{1}\right)\left(d^{\frac{-1}{2}}\lambda_{N}^{\frac{1}{2N}-1}\right)}{4N\sqrt{2}\beta_{i}}\frac{\alpha^{\prime}}{\tau^{3}}$
$-\frac{-\left(38N-11\right)\lambda_{N-1}^{2}b_{1}+48N\lambda_{N}\lambda_{N-1}b_{2}+48N^{2}\lambda_{N}^{2}b_{2}-48N^{2}\lambda_{N}^{2}b_{3}+76N\lambda_{N-2}\lambda_{N}b_{1}}{N^{2}2^{5}\sqrt{2}\beta_{i}}\left(d^{\frac{-3}{2}}\lambda_{N}^{\frac{1}{2N}-2}\right)\frac{\alpha^{\prime 2}}{\tau^{5}}$
$+\left[-\left(1262N^{2}-5133N+6560\right)\lambda_{N-1}^{3}b_{1}+4N^{2}\left(82N-25\right)\lambda_{N-2}\lambda_{N}\lambda_{N-1}b_{1}-328N^{3}\lambda_{N-3}\lambda_{N}^{2}b_{1}\right$.
$+8N^{2}\left(22N-7\right)\lambda_{N}\lambda_{N-1}^{2}b_{2}+112N^{3}\lambda_{N}^{2}\lambda_{N-1}b_{2}+128N^{4}\lambda_{N}^{3}b_{2}-352N^{3}\lambda_{N-2}\lambda_{N}^{2}b_{2}$
$\left.-112N^{3}\lambda_{N}^{2}\lambda_{N-1}b_{3}+128N^{4}\lambda_{N}^{3}b_{4}\right]\frac{d^{\frac{-5}{2}}\lambda_{N}^{\frac{1}{2N}-3}}{128N^{4}\sqrt{2}\beta_{i}}\frac{\alpha^{\prime 3}}{\tau^{7}}+\mathcal{O}\left(\frac{1}{\tau^{9}}\right).$
To compare this result with perturbative solution (32) as $\alpha^{\prime}\rightarrow 0$, we can fix $\gamma=\frac{1}{\sqrt{\alpha^{\prime}}}$ and the coefficients $\lambda_{k}$ and $b_{k}$:
$\lambda_{N}$ $=$ $\left(\frac{1}{2d}\right)^{N},$
$\lambda_{N-1}$ $=$ $dN\left(\frac{1}{2d}\right)^{N}\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4},$
$\lambda_{N-2}$ $=$ $\left(\frac{1}{2}\left(N-1\right)\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)^{2}-\frac{8}{9}\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{6}\right)d^{2}N\left(\frac{1}{2d}\right)^{N},$
$\lambda_{N-3}$ $=$ $\left[\left(\frac{64}{45}-\frac{8}{9}N\right)\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{6}\mathrel{{\mathop{\sum}\limits_{k=1}^{d}}}\beta_{k}^{4}+\frac{8}{5}\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{8}+\frac{N^{2}-3N+2}{6}\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)^{3}\right$.
$\left.+\frac{8192}{5}\left(\mathrel{{\mathop{\sum}\limits_{k=1}^{d}}}\beta_{k}^{4}\right)^{2}c_{4,1}+\frac{4096}{5}\left(\mathrel{{\mathop{\sum}\limits_{k=1}^{d}}}\beta_{k}^{8}\right)c_{4,0}\right]d^{3}N\left(\frac{1}{2d}\right)^{N},$
$\ldots$ (36)
and
$b_{1}\left(\beta_{i}\right)$ $=$ $2\beta_{i}^{2},$
$b_{2}\left(\beta_{i}\right)$ $=$ $-d\beta_{i}^{2}\left(\beta_{i}^{2}-\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right),$
$b_{3}\left(\beta_{i}\right)$ $=$ $\frac{d\beta_{i}^{2}}{3}\left(3\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{k}^{4}-8d\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{6}+8d\beta_{i}^{4}-3\beta_{i}^{2}\right),$
$b_{4}\left(\beta_{i}\right)$ $=$ $\frac{d\beta_{i}^{2}}{9}\left[8d^{2}\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{6}\right)\left(\mathrel{{\mathop{\sum}\limits_{k=1}^{d}}}\beta_{k}^{4}\right)+72d^{2}\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{8}\right)-9\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)\right$.
$+\beta_{i}^{2}\left(16d^{2}\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{6}\right)+9\right)-24d^{2}\beta_{i}^{4}\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)-72d^{2}\beta_{i}^{6}$
$\left.+73728d^{2}c_{4,1}\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)\left(\left(\mathrel{{\mathop{\sum}\limits_{k=1}^{d}}}\beta_{k}^{4}\right)-\beta_{i}^{2}\right)+36864d^{2}c_{4,0}\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{8}-\beta_{i}^{6}\right)\right],$
$\ldots,$ (37)
where
$\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}b_{1}\left(\beta_{i}\right)=2,\qquad\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}b_{k}\left(\beta_{i}\right)=0,\qquad k>1,$ (38)
Now, we study the singularity structure of the non-perturbative solution (33) through the associated Kretschmann scalar:
$R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ $=$ $2\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}^{2}\right)^{2}+2\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}^{4}\right)+8\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}^{2}\dot{H}_{i}\right)+4\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\dot{H}_{i}^{2}\right)$ (39)
$=$ $2\left[\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\frac{1}{\beta_{i}^{2}}\right)^{2}+\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\frac{1}{\beta_{i}^{4}}\right]\frac{1}{e^{4\Phi\left(\tau\right)}}\left(\sqrt{\alpha^{\prime}}\ddot{\Phi}A\left(\sigma\right)-\sqrt{2d}\dot{\Phi}B\left(\sigma\right)\right)^{4}$
$+8\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\frac{1}{\beta_{i}^{3}}\frac{1}{e^{3\Phi\left(\tau\right)}}\left(\sqrt{\alpha^{\prime}}\ddot{\Phi}A\left(\sigma\right)-\sqrt{2d}\dot{\Phi}B\left(\sigma\right)\right)^{2}$
$\times\left[\sqrt{\alpha^{\prime}}\dddot{\Phi}A\left(\sigma\right)-\sqrt{\alpha^{\prime}}\ddot{\Phi}\dot{\Phi}A\left(\sigma\right)+\sqrt{\alpha^{\prime}}\ddot{\Phi}\frac{d}{d\tau}A\left(\sigma\right)\right$.
$\left.-\sqrt{2d}\ddot{\Phi}B\left(\sigma\right)+\sqrt{2d}\dot{\Phi}^{2}B\left(\sigma\right)-\sqrt{2d}\dot{\Phi}\frac{d}{d\tau}B\left(\sigma\right)\right]$
$+4\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\frac{1}{\beta_{i}^{2}}\frac{1}{e^{2\Phi\left(\tau\right)}}\left[\sqrt{\alpha^{\prime}}\dddot{\Phi}A\left(\sigma\right)-\sqrt{\alpha^{\prime}}\ddot{\Phi}\dot{\Phi}A\left(\sigma\right)+\sqrt{\alpha^{\prime}}\ddot{\Phi}\frac{d}{d\tau}A\left(\sigma\right)\right$.
$\left.-\sqrt{2d}\ddot{\Phi}B\left(\sigma\right)+\sqrt{2d}\dot{\Phi}^{2}B\left(\sigma\right)-\sqrt{2d}\dot{\Phi}\frac{d}{d\tau}B\left(\sigma\right)\right]^{2}.$
For simplicity, we define
$A\left(\sigma\right)\equiv\mathrel{{\mathop{\sum}\limits_{k=1}^{\infty}}}\frac{b_{k}}{1+\sigma^{2k-2}},\;B\left(\sigma\right)\equiv\mathrel{{\mathop{\sum}\limits_{k=1}^{\infty}}}\frac{k\sigma^{2k-1}b_{k+1}}{\left(1+\sigma^{2k}\right)^{2}}.$ (40)
To analyze the singularities of the Kretschmann scalar, we introduce
$S\left(\sigma\right)\equiv\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k},$ (41)
in terms of which the scalar curvature invariant takes the compact form
$R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}=\left(\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}\right)^{-\varrho}F\left(\tau\right),$ (42)
where $\varrho\equiv 8-\frac{2}{N}$ and
$F\left(\tau\right)$ (43)
$=$ $\frac{2}{\left(2N\right)^{4}}\left(\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\frac{1}{\beta_{i}^{2}}\right)^{2}+\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\frac{1}{\beta_{i}^{4}}\right)\left(-\sqrt{\alpha^{\prime}}\left[S\left(\tau\right)\frac{d^{2}}{d\tau^{2}}S\left(\tau\right)-\left(\frac{d}{d\tau}S\left(\tau\right)\right)^{2}\right]A\left(\sigma\right)+\sqrt{2d}B\left(\sigma\right)S\left(\tau\right)\frac{d}{d\tau}S\left(\tau\right)\right)^{4}$
$+8\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\frac{1}{\beta_{i}^{3}}\frac{1}{\left(2N\right)^{4}}S\left(\tau\right)^{1-\frac{1}{2N}}\left(-\sqrt{\alpha^{\prime}}\left[S\left(\tau\right)\frac{d^{2}}{d\tau^{2}}S\left(\tau\right)-\left(\frac{d}{d\tau}S\left(\tau\right)\right)^{2}\right]A\left(\sigma\right)+\sqrt{2d}B\left(\sigma\right)S\left(\tau\right)\frac{d}{d\tau}S\left(\tau\right)\right)^{2}$
$\times\left[-\sqrt{\alpha^{\prime}}\left[2NS\left(\tau\right)^{2}\frac{d^{3}}{d\tau^{3}}S\left(\tau\right)-\left(6N-1\right)S\left(\tau\right)\frac{d}{d\tau}S\left(\tau\right)\frac{d^{2}}{d\tau^{2}}S\left(\tau\right)+\left(4N-1\right)\left(\frac{d}{d\tau}S\left(\tau\right)\right)^{3}\right]A\left(\sigma\right)\right$.
$-2N\sqrt{\alpha^{\prime}}\left[S\left(\tau\right)^{2}\frac{d^{2}}{d\tau^{2}}S\left(\tau\right)-S\left(\tau\right)\left(\frac{d}{d\tau}S\left(\tau\right)\right)^{2}\right]\frac{d}{d\tau}A\left(\sigma\right)$
$\left.+\sqrt{2d}\left[2NS\left(\tau\right)^{2}\frac{d^{2}}{d\tau^{2}}S\left(\tau\right)-\left(2N-1\right)S\left(\tau\right)\left(\frac{d}{d\tau}S\left(\tau\right)\right)^{2}\right]B\left(\sigma\right)+2N\sqrt{2d}S\left(\tau\right)^{2}\frac{d}{d\tau}S\left(\tau\right)\frac{d}{d\tau}B\left(\sigma\right)\right]$
$+4\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\frac{1}{\beta_{i}^{2}}\frac{1}{\left(2N\right)^{4}}S\left(\tau\right)^{2-\frac{1}{N}}$
$\times\left[-\sqrt{\alpha^{\prime}}\left[2NS\left(\tau\right)^{2}\frac{d^{3}}{d\tau^{3}}S\left(\tau\right)-\left(6N-1\right)S\left(\tau\right)\frac{d}{d\tau}S\left(\tau\right)\frac{d^{2}}{d\tau^{2}}S\left(\tau\right)+\left(4N-1\right)\left(\frac{d}{d\tau}S\left(\tau\right)\right)^{3}\right]A\left(\sigma\right)\right$.
$-2N\sqrt{\alpha^{\prime}}\left[S\left(\tau\right)^{2}\frac{d^{2}}{d\tau^{2}}S\left(\tau\right)-S\left(\tau\right)\left(\frac{d}{d\tau}S\left(\tau\right)\right)^{2}\right]\frac{d}{d\tau}A\left(\sigma\right)$
$\left.+\sqrt{2d}\left[2NS\left(\tau\right)^{2}\frac{d^{2}}{d\tau^{2}}S\left(\tau\right)-\left(2N-1\right)S\left(\tau\right)\left(\frac{d}{d\tau}S\left(\tau\right)\right)^{2}\right]B\left(\sigma\right)+2N\sqrt{2d}S\left(\tau\right)^{2}\frac{d}{d\tau}S\left(\tau\right)\frac{d}{d\tau}B\left(\sigma\right)\right]^{2}.$
Using (40) and (41), one finds
$\frac{d}{d\tau}S\left(\sigma\right)$ $=$ $\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}2k\lambda_{k}\sigma^{2k-1}\sqrt{\frac{2d}{\alpha^{\prime}}},$
$\frac{d^{2}}{d\tau^{2}}S\left(\sigma\right)$ $=$ $\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}2k\left(2k-1\right)\lambda_{k}\sigma^{2k-2}\frac{2d}{\alpha^{\prime}},$
$\frac{d^{3}}{d\tau^{3}}S\left(\sigma\right)$ $=$ $\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}2k\left(2k-1\right)\left(2k-2\right)\lambda_{k}\sigma^{2k-3}\left(\frac{2d}{\alpha^{\prime}}\right)^{\frac{3}{2}},$
$\frac{d}{d\tau}A\left(\sigma\right)$ $=$ $-\mathrel{{\mathop{\sum}\limits_{k=1}^{\infty}}}\frac{\left(2k-2\right)b_{k}\sigma^{2k-3}}{\left(1+\sigma^{2k-2}\right)^{2}},$
$\frac{d}{d\tau}B\left(\sigma\right)$ $=$ $\mathrel{{\mathop{\sum}\limits_{k=1}^{\infty}}}\frac{k\left(2k-1\right)\sigma^{2k-2}b_{k+1}\left(1+\sigma^{2k}\right)^{2}\sqrt{\frac{2d}{\alpha^{\prime}}}-\left(2k\right)^{2}\sigma^{2k-1}b_{k+1}\left(1+\sigma^{2k}\right)\sigma^{2k-1}\sqrt{\frac{2d}{\alpha^{\prime}}}}{\left(1+\sigma^{2k}\right)^{4}}.$
It is evident that (40), (41) and (LABEL:eq:dSAB) are free from singularities, since the presence of $1+\sigma^{2n}$ (with even powers of $\sigma$) in the denominator never vanishes. Therefore, $F\left(\tau\right)$, which is constructed from finite combinations of these quantities through addition, subtraction, and multiplication, is a regular function. Consequently, the spacetime and dilaton fields become singular if and only if
$S\left(\sigma\right)=\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}=0,$ (45)
admits real roots.
.3 Non-perturbative solutions which cover $\alpha^{\prime 3}$ as $\alpha^{\prime}\rightarrow 0$ (special example):
To cover the perturbative solution (32) up to the order of $\alpha^{\prime 3}$ as $\alpha^{\prime}\rightarrow 0$, and including the known coefficients $c_{1,0}=-\frac{1}{8}$, $c_{2,0}=\frac{1}{64}$, $c_{3,0}=-\frac{1}{3.2^{7}}$, $c_{4,0}=\frac{1}{2^{12}}-\frac{3}{2^{12}}\zeta\left(3\right)$, $c_{4,1}=\frac{1}{2^{16}}+\frac{1}{2^{12}}\zeta\left(3\right)$, we need to choose $N=3$ and $d=3$, which denotes the singularity regions of the four-dimensional Schwarzschild and rotating, accreting black holes. The corresponding non-perturbative solution (33) is given by:
$\Phi\left(\tau\right)$ $=$ $-\frac{1}{6}\log\left(\lambda_{0}+\lambda_{1}\sigma^{2}+\lambda_{2}\sigma^{4}+\lambda_{3}\sigma^{6}\right),\qquad\sigma^{2}\equiv 6\frac{\tau^{2}}{\alpha^{\prime}},$
$f_{i}\left(H_{i}\right)$ $=$ $-\frac{2\beta_{i}}{\sqrt{\alpha^{\prime}}}\left(\lambda_{0}+\lambda_{1}\sigma^{2}+\lambda_{2}\sigma^{4}+\lambda_{3}\sigma^{6}\right)^{-\frac{1}{6}},$
$g_{i}\left(H_{i}\right)$ $=$ $-\frac{\left(\lambda_{1}\sigma+2\lambda_{2}\sigma^{3}+3\lambda_{3}\sigma^{5}\right)^{2}\left(\frac{b_{1}}{2}+\frac{b_{2}}{1+\sigma^{2}}+\frac{b_{3}}{1+\sigma^{4}}+\frac{b_{4}}{1+\sigma^{6}}\right)}{3\alpha^{\prime}\left(\lambda_{0}+\lambda_{1}\sigma^{2}+\lambda_{2}\sigma^{4}+\lambda_{3}\sigma^{6}\right)^{2}},$
$H_{i}\left(\tau\right)$ $=$ $\frac{1}{\sqrt{\alpha^{\prime}}\beta_{i}}\left(\lambda_{0}+\lambda_{1}\sigma^{2}+\lambda_{2}\sigma^{4}+\lambda_{3}\sigma^{6}\right)^{-\frac{11}{6}}\left[-\left(\lambda_{0}+\lambda_{1}\sigma^{2}+\lambda_{2}\sigma^{4}+\lambda_{3}\sigma^{6}\right)\times\right$. (46)
$\left(2\lambda_{1}+12\lambda_{2}\sigma^{2}+30\lambda_{3}\sigma^{3}\right)\left(\frac{b_{1}}{2}+\frac{b_{2}}{1+\sigma^{2}}+\frac{b_{3}}{1+\sigma^{4}}+\frac{b_{4}}{1+\sigma^{6}}\right)$
$+\left(2\lambda_{1}\sigma+4\lambda_{2}\sigma^{3}+6\lambda_{3}\sigma^{5}\right)\left(\lambda_{0}+\lambda_{1}\sigma^{2}+\lambda_{2}\sigma^{4}+\lambda_{3}\sigma^{6}\right)\times$
$\left(\frac{\sigma b_{2}}{\left(1+\sigma^{2}\right)^{2}}+\frac{2\sigma^{3}b_{3}}{\left(1+\sigma^{4}\right)^{2}}+\frac{3\sigma^{5}b_{4}}{\left(1+\sigma^{6}\right)^{2}}\right)+\left(2\lambda_{1}\sigma+4\lambda_{2}\sigma^{3}+6\lambda_{3}\sigma^{5}\right)^{2}\times$
$\left.\left(\frac{b_{1}}{2}+\frac{b_{2}}{1+\sigma^{2}}+\frac{b_{3}}{1+\sigma^{4}}+\frac{b_{4}}{1+\sigma^{6}}\right)\right],$
with $\lambda_{3}=\frac{1}{216}$, $\lambda_{2}=\frac{1}{216}$, $\lambda_{1}=\frac{19}{17496}$, $\text{$\lambda_{0}$}=\frac{6303744c_{4,0}+26763264c_{4,1}+5701}{262440}$ and
$b_{1}\left(\beta_{i}\right)$ $=$ $2\beta_{i}^{2},$
$b_{2}\left(\beta_{i}\right)$ $=$ $3\beta_{i}^{2}\left(\frac{11}{27}-\beta_{i}^{2}\right),$
$b_{3}\left(\beta_{i}\right)$ $=$ $\beta_{i}^{2}\left(24\beta_{i}^{4}-3\beta_{i}^{2}-\frac{245}{81}\right),$
$b_{4}\left(\beta_{i}\right)$ $=$ $\frac{\beta_{i}^{2}}{3}\left(270336c_{41}\left(\frac{11}{27}-\beta_{i}^{2}\right)+331776c_{40}\left(\frac{19}{243}-\beta_{i}^{6}\right)-648\beta_{i}^{6}-88\beta_{i}^{4}+\frac{931\beta_{i}^{2}}{27}+\frac{38047}{729}\right),$ (47)
where $c_{4,0}=\frac{1}{2^{12}}-\frac{3}{2^{12}}\zeta\left(3\right)$ and $c_{4,1}=\frac{1}{2^{16}}+\frac{1}{2^{12}}\zeta\left(3\right)$. It is worth noting that
$\mathrel{{\mathop{\sum}\limits_{i=1}^{3}}}b_{1}\left(\beta_{i}\right)=2,$ (48)
and
$\mathrel{{\mathop{\sum}\limits_{i=1}^{3}}}b_{2}\left(\beta_{i}\right)=\mathrel{{\mathop{\sum}\limits_{i=1}^{3}}}b_{3}\left(\beta_{i}\right)=\mathrel{{\mathop{\sum}\limits_{i=1}^{3}}}b_{4}\left(\beta_{i}\right)=0.$ (49)
The corresponding Kretschmann scalar is:
$R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ $=$ $2\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}^{2}\right)^{2}+2\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}^{4}\right)+8\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}^{2}\dot{H}_{i}\right)+4\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\dot{H}_{i}^{2}\right)$ (50)
$=$ $\left(\left(15336\zeta\left(3\right)+61187\right)+13680\sigma^{2}+1283040\sigma^{4}+2099520\sigma^{6}\right)^{-\varrho}F\left(\tau\right),$
Then, it is easy to check that this equation has no real root, implying the singularities of four-dimensional Schwarzschild and rotating, accreting black holes are removed. Finally, we wish to note that although the coefficients $c_{k\geq 5}$ are unknown, in principle, they can be determined perturbatively by the worldsheet anomaly cancellation.
.4 Regularity of the string coupling and curvature invariant in the Einstein frame
To examine the regularity of the string coupling $g_{s}=\exp\left(\phi\right)$ , we first recall the definition of the physical dilaton,
$2\phi$ $=$ $\Phi+\log\sqrt{\left|g_{ij}\right|}$ (51)
$=$ $-\frac{1}{2N}\log\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}+\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\int H_{i}d\tau,$
where the Hubble parameter is given by
$H_{i}=\frac{S\left(\tau\right)^{\frac{1}{2N}-2}}{2N\beta_{i}}\left(-\sqrt{\alpha^{\prime}}A\left(\sigma\right)\left(S\left(\tau\right)\frac{d^{2}}{d\tau^{2}}S\left(\tau\right)-\left(\frac{d}{d\tau}S\left(\tau\right)\right)^{2}\right)+\sqrt{2d}B\left(\sigma\right)S\left(\tau\right)\frac{d}{d\tau}S\left(\tau\right)\right),$ (52)
with $A\left(\sigma\right)$ and $B\left(\sigma\right)$ defined in (40), and $S\left(\tau\right)$ given by (41). To verify the regularity of the physical dilaton, we need to check whether
$S\left(\tau\right)=\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k},$ (53)
vanishes. If $S\left(\tau\right)\neq 0$, then the first term in the expression (51) for $\phi$ remains regular. Furthermore, from the coordinate transformation between the inner horizon of the Schwarzschild black hole and the Kasner universe, the region where the Kasner description is valid is bounded by
$0<\tau\ll\tau_{max}.$ (54)
For instance, in the four-dimensional case, the maximal value of $\tau$ is $\tau_{max}=\frac{4}{3}GM$. In this integration domain, the integrand $H_{i}$ of $\int H_{i}d\tau$ is both continuous and bounded, implying that the integral is convergent. Consequently, the condition $S\left(\sigma\right)=\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}\neq 0$ guarantees that $\phi$ is regular, and hence the string coupling $g_{s}$ remains finite. Therefore, loop corrections need not be considered in our analysis.
On the other hand, to study the regularity of the curvature invariant $R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ in the Einstein frame, we recall the relation between the string frame and Einstein frame metrics:
$g_{\mu\nu}^{E}\hskip 0.0pt=e^{-\frac{4\hskip 0.0pt\left(\phi-\phi_{0}\right)\hskip 0.0pt}{d-1}}g_{\mu\nu}.$ (55)
The corresponding Hubble parameter in the Einstein frame is related to that in the string frame by
$H_{i}^{E}=e^{\frac{2\phi}{d-1}}\left(H_{i}-\frac{2\dot{\phi}}{d-1}\right)$ (56)
Based on this relation, the curvature invariant in the Einstein frame can be written as
$\left(R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}\right)^{E}$ (57)
$=$ $2\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\left(H_{i}^{E}\right)^{2}\right)^{2}+2\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\left(H_{i}^{E}\right)^{4}\right)+8\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\left(H_{i}^{E}\right)^{2}\frac{d}{dt}H_{i}^{E}\right)+4\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\left(\frac{d}{dt}H_{i}^{E}\right)^{2}\right)$
$=$ $2e^{\frac{8\phi}{d-1}}\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\left(H_{i}-\frac{1}{d-1}\left(\dot{\Phi}+\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}H_{j}\right)\right)^{2}\right)^{2}+2e^{\frac{8\phi}{d-1}}\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\left(H_{i}-\frac{1}{d-1}\left(\dot{\Phi}+\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}H_{j}\right)\right)^{4}\right)$
$+8e^{\frac{8\phi}{d-1}}\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\left(H_{i}-\frac{1}{d-1}\left(\dot{\Phi}+\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}H_{j}\right)\right)^{2}\left[\frac{H_{i}}{d-1}\left(\dot{\Phi}+\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}H_{j}\right)\right$.
$\left.-\frac{1}{\left(d-1\right)^{2}}\left(\dot{\Phi}+\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}H_{j}\right)^{2}+\dot{H}_{i}-\frac{1}{d-1}\left(\ddot{\Phi}+\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\dot{H}_{j}\right)\right]$
$+4e^{\frac{8\phi}{d-1}}\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\left(\frac{H_{i}}{d-1}\left(\dot{\Phi}+\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}H_{j}\right)-\frac{1}{\left(d-1\right)^{2}}\left(\dot{\Phi}+\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}H_{j}\right)^{2}+\dot{H}_{i}-\frac{1}{d-1}\left(\ddot{\Phi}+\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\dot{H}_{j}\right)\right)^{2}.$
The quantities involved are
$H_{i}$ $=$ $\frac{S\left(\tau\right)^{\frac{1}{2N}}}{\beta_{i}}\left(\sqrt{\alpha^{\prime}}\ddot{\Phi}A\left(\sigma\right)-\sqrt{2d}\dot{\Phi}B\left(\sigma\right)\right),$
$\dot{H}_{i}$ $=$ $S\left(\tau\right)^{\frac{1}{2N}}\left(\frac{\sqrt{\alpha^{\prime}}\dddot{\Phi}}{\beta_{i}}A\left(\sigma\right)-\frac{\sqrt{\alpha^{\prime}}\ddot{\Phi}\dot{\Phi}}{\beta_{i}}A\left(\sigma\right)+\frac{\sqrt{\alpha^{\prime}}\ddot{\Phi}}{\beta_{i}}\frac{d}{d\tau}A\left(\sigma\right)\right$. (58)
$\left.-\frac{\sqrt{2d}\ddot{\Phi}}{\beta_{i}}B\left(\sigma\right)+\frac{\sqrt{2d}\dot{\Phi}^{2}}{\beta_{i}}B\left(\sigma\right)-\frac{\sqrt{2d}\dot{\Phi}}{\beta_{i}}\frac{d}{d\tau}B\left(\sigma\right)\right),$
and
$\dot{\Phi}$ $=$ $-\frac{\frac{d}{d\tau}S\left(\tau\right)}{2NS\left(\tau\right)},$
$\ddot{\Phi}$ $=$ $-\frac{S\left(\tau\right)\frac{d^{2}}{d\tau^{2}}S\left(\tau\right)-\left(\frac{d}{d\tau}S\left(\tau\right)\right)^{2}}{2NS\left(\tau\right)^{2}},$
$\dddot{\Phi}$ $=$ $-\frac{S\left(\tau\right)^{2}\frac{d^{3}}{d\tau^{3}}S\left(\tau\right)-3S\left(\tau\right)\frac{d}{d\tau}S\left(\tau\right)\frac{d^{2}}{d\tau^{2}}S\left(\tau\right)+2\left(\frac{d}{d\tau}S\left(\tau\right)\right)^{3}}{2NS\left(\tau\right)^{3}}.$ (59)
Following the same analysis as in Appendix B, we find that the regularity of all quantities depends on whether $S\left(\tau\right)$ vanishes. If $S\left(\tau\right)=\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}\neq 0$, then all terms $\phi$, $H_{i},\dot{H}_{i},\dot{\Phi},\ddot{\Phi}$ and $\dddot{\Phi}$ are regular. Consequently, the curvature invariant $\left(R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}\right)^{E}$ is also free of singularities in the Einstein frame.